### v2

# Non-universal surface magnetoelectric response in antiferromagnetic topological insulators

Abstract
The electronic ground state of a three-dimensional (3D) band insulator with time-reversal ($\Theta$) symmetry or time-reversal times a discrete translation ($\Theta T_{1/2}$) symmetry is classified by a $\mathbb{Z}_{2}$-valued topological invariant and characterized by quantized magnetoelectric response. Here we demonstrate by explicit calculation in model $\mathbb{Z}_{2}$ topological insulator thin-films that whereas the magnetoelectric response is localized at the surface in the $\Theta$ symmetry (non-magnetic) case, it is non-universally partitioned between surface and interior contributions in the $\Theta T_{1/2}$ (anti-ferromagnetic) case, while remaining quantized. Within our model the magnetic field induced polarization arises entirely from an anomalous ${\cal N}=0$ Landau level subspace within which the projected Hamiltonian is a generalized Su-Schrieffer-Heeger model whose topological properties are consistent with those of the starting 3D model. We identify a novel connection between the ground state geometry of that 3D model and surface-interior-partitioning in thin films.
Introduction— The realization that material properties can depend qualitatively [1, 2] on topological data encoded in the crystal-momentum-dependence of its occupied electronic eigenspace projectors has been an important achievement in modern condensed matter physics. The physical import of this mathematical structure was first recognized in theoretical studies [3, 4] of the integer quantum Hall effect (IQHE) in which the Hall conductivity of two-dimensional band insulators was identified as the (quantized) Chern number of the vector bundle of occupied (electronic) Bloch states over the Brillouin zone (BZ) torus [5]. Subsequently other, more subtle, topological invariants of band insulating ground states have been found to be physically relevant. Among these, three-dimensional (3D) band insulators with time-reversal symmetry (TRS, denoted as $\Theta$) or generalized TRS ($\Theta T_{1/2}$), which have a quantized magnetoelectric response [6, 7, 8],
$\alpha^{il}=\left.\frac{\partial P^{i}}{\partial B^{l}}\right|_{\begin{subarray}{c}\bm{E}=\bm{0}\\ \bm{B}=\bm{0}\end{subarray}}=\left.\frac{\partial M^{l}}{\partial E^{i}}\right|_{\begin{subarray}{c}\bm{E}=\bm{0}\\ \bm{B}=\bm{0}\end{subarray}}=\delta^{il}\alpha,$ (1)
have been a prominent focus. In Eq. (1) superscript indices identify Cartesian components, ${\mathbf{P}}$ and ${\mathbf{M}}$ are the (electronic) polarization and orbital magnetization, and ${\mathbf{E}}$ and ${\mathbf{B}}$ are uniform dc electric and magnetic fields. In both non-magnetic ($\Theta$ symmetric) and anti-ferromagnetic (AFM) ($\Theta T_{1/2}$ symmetric [9]) cases, $\alpha=0$ or $e^{2}/2hc$ (mod $e^{2}/hc$) and is determined, in part, by the $\mathbb{Z}_{2}$ classification of the ground state [6, 7, 8]. Indeed, an explicit expression has been proposed [6] that relates $\alpha$ to the bulk BZ integral of the Chern-Simons 3-form in the space of occupied Bloch states. The value of $\alpha$ implied by this bulk expression ($\alpha_{\text{CS}}$), quoted explicitly below, is determined modulo $e^{2}/hc$ by the $\mathbb{Z}_{2}$ index of the vector bundle of occupied Bloch states [6, 9, 10, 11].
Figure 1: Landau levels (LLs). (a) Schematic topological insulator thin film energy spectrum in the absence (black solid) and presence (LL spectra) of a uniform dc magnetic field. The low-energy Hamiltonian for an AFM magnetic configuration acts in the ${\cal N}=0$ LL subspace as a generalized 1D SSH model (see main text) with four electronic states (represented by colored circles) per unit cell (green rectangle), one state at each surface of two magnetic layers (dashed ovals) with opposite magnetization (represented by red and blue colors). The states at each surface of every (interior) layer are coupled to the static magnetic moments in two layers that have opposite spin-orientations, one in the same layer with exchange coupling $J_{\text{S}}$ and one in an adjacent layer with exchange coupling $J_{\text{D}}$. Thus, the effective Zeeman energies ±$\pm(J_{\text{S}}-J_{\text{D}})$ (represented by light and dark green arrows) alternate in sign across the layers. Schematic charge distributions of anomalous ${\cal N}=0$ LLs (red curves) and ${\cal N}\neq 0$ LLs (purple dashed curves) in TI thin films: (b) with magnetic dopants only in the exterior layers and (c) with interior magnetic moments and A-type antiferromagnetism.
The relationship between $\alpha_{\text{CS}}$ and the physical magnetoelectric response ($\alpha_{\text{me}}$) of finite volume $\mathbb{Z}_{2}$ topological insulators (TIs) with surfaces is non-trivial [8, 12, 13, 14, 15]. In non-magnetic thin films with surface magnetic dopants that have nonzero surface-normal magnetization components that is opposite on top and bottom surfaces, $\alpha_{\text{me}}$ approaches $\alpha_{\text{CS}}$ in the thick film limit. Here we demonstrate by explicit calculation that whereas the quantized magnetoelectric response is localized at the surface in the non-magnetic case, it is non-universally partitioned between surface and interior contributions in the AFM case. In both cases the response of the (electronic) charge density to magnetic field is local in the sense that it occurs only in the spatial vicinity of static magnetic moments that break $\Theta$ symmetry. The difference is that these are present throughout the film interior in the AFM case. The dipole moment of the charge density that is induced by the magnetic field in the interior region of an AFM film is non-universal, but for sufficiently thick films we find that the sum of interior and surface contributions to the dipole moment is quantized (mod $e^{2}/hc$) at either $\alpha_{\text{me}}=0$ or $e^{2}/2hc$ depending on the $\mathbb{Z}_{2}$ classification of the bulk ground state; in AFM (non-magnetic) thin films (with magnetic surface dopants) $\Theta T_{1/2}$ ($\Theta$) symmetry is broken and therefore, e.g., $\alpha_{\text{me}}=0$ cannot be protected by that symmetry, instead it is protected by the bulk $\Theta T_{1/2}$ ($\Theta$)-symmetry-induced topological classification of the ground state. This property, that the sum of 3D interior and surface contributions is in accord with 3D bulk topological constraints even when the separate contributions are non-quantized and non-universal, is reminiscent of the frequently discussed non-universal Hall current partition between interior and edge [16, 17, 18, 19] in the IQHE.
Our explicit calculations employ a model that can describe both non-magnetic and AFM layered materials, and for each case can support ground states characterized by a trivial or non-trivial $\mathbb{Z}_{2}$ index. The model provides a realistic description of A-type AFM TIs, similar to the paradigmatic AFM TI MnBi2Te4 [20, 21, 22, 23, 24, 25, 26, 27, 28, 29], but is simplified in a way that is convenient for explicit inclusion of magnetic fields that are parallel to the stacking axis [30]. We use a continuum version of this model to calculate electronic charge distributions across thin films and a lattice regularized version to address the topological magnetoelectric response of bulk crystals. As illustrated in Fig. 1 (a), our thin film calculations show that the magnetic field induced dipole moment arises entirely from the model’s anomalous $\mathcal{N}=0$ Landau level (LL) subspace, the dynamics of which are governed by a generalized Su-Schrieffer-Heeger (SSH) Hamiltonian. Because the $\mathcal{N}\neq 0$ LL eigenfunctions are symmetric across the thin film, as illustrated in Fig. 1 (b) and (c), they do not contribute to the magnetic field induced dipole moment. The absence of a dipole moment in the interior of a non-magnetic film is illustrated schematically in Fig. 1 (b), and its presence in an AFM film in Fig. 1 (c).
Anomalous Landau levels—An effective low-energy Hamiltonian for the electronic states in the layered materials of interest can be constructed [31, 32, 30] using Dirac cones, one associated with top and bottom surface of each layer, that are coupled by hybridizing states both within and between layers:
$\hat{H}=$ $\sum_{{\mathbf{k}}_{\perp},ij,\mu\gamma}\Big{[}\Big{(}\,(-1)^{i}\hbar v_{D}(\hat{z}\times{\mathbf{\sigma}})\cdot{\mathbf{k}}_{\perp}+m_{i}\sigma_{z}\Big{)}_{\mu,\gamma}\delta_{ij}$
$\qquad\qquad+\Delta_{ij}(1-\delta_{ij})\delta_{\mu,\gamma}\Big{]}\hat{c}_{{\mathbf{k}}_{\perp},i,\mu}^{\dagger}\hat{c}_{{\mathbf{k}}_{\perp},j,\gamma}.$ (2)
Here we have chosen $\hat{\bm{z}}$ as the stacking axis, ${\mathbf{k}}_{\perp}=(k_{x},k_{y})$, $i,j\in\{0,\ldots,2N-1\}$ where $i=2(l-1)$ $(i=2l-1)$ labels states that are associated with the bottom (top) surface of layer $l\in\{1,\ldots,N\}$, $\mu,\gamma\in\{\uparrow,\downarrow\}$ are spin labels, $\Delta_{ij}$ are the intra- and inter-layer hybridization parameters, $v_{{}_{D}}$ is the velocity of an isolated Dirac cone, and $m_{i}$ are mass terms that account for exchange coupling to static local magnetic moments (which we take to be aligned with the stacking direction). We will consider only even layer number $N$ so that AFM thin films do not have uncompensated moments, and in our model $m_{i}$ is a mass term at surface $i$ that accounts for exchange coupling to static local magnetic moments. We set $m_{i}=\sum_{l}J_{il}M_{l}$ with dimensionless magnetization $M_{l}=0$ if layer $l$ is non-magnetic and ±$\pm 1$ if layer $l$ is magnetic (for antiferromagnetic configuration we take $M_{l}=(-1)^{l}$); the Dirac-like states at the top or bottom surface of layer $l$ are coupled to the static magnetization of that layer (via $J_{\text{S}}$) and to that of the adjacent layer nearest that surface (via $J_{\text{D}}$), and hybridized only to adjacent surfaces with hopping $\Delta_{\text{S}}$ for opposite surfaces of the same layer and $\Delta_{\text{D}}$ for the closest surface of the adjacent layer (see Fig. 1 of Ref. [30]). In thin film models the Dirac-like states of the outermost surfaces do not have $J_{\text{D}}$ or $\Delta_{\text{D}}$ coupling.
We account for an external magnetic field ${\mathbf{B}}=-B\hat{{\mathbf{z}}}$ though the usual prescription  [^1] See, e.g., Chapter 2 of Winkler [56]. of mapping a ${\mathbf{k}}\cdot{\mathbf{p}}$ model to an envelope function approximated (real-space) Hamiltonian that is then minimally-coupled to the electromagnetic field: $\hbar{\mathbf{k}}\rightarrow-i\hbar\bm{\nabla}-(eBx/c)\hat{{\mathbf{y}}}$ in the Landau gauge. Here $e>0$ is the elementary charge unit. Applying this recipe to Eq. (S1) and seeking energy eigenfunctions of the form $\Psi_{E,q_{y}}(x,y)=e^{iq_{y}y}\Phi_{E,q_{y}}(x)/\sqrt{L_{y}}$, where $\Phi_{E,q_{y}}(x)$ has $2N\times 2$ components that correspond with layer and spin label degrees of freedom, results in the individual Dirac Hamiltonian $H_{D}$ of each surface being recast [34] in terms of the LL raising and lowering (differential) operators $a^{\dagger}\equiv(\tilde{x}-\partial_{\tilde{x}})/\sqrt{2}$ and $a\equiv(\tilde{x}+\partial_{\tilde{x}})/\sqrt{2}$,
$H_{D}=\hbar v_{{}_{D}}(\hat{z}\times{\mathbf{\sigma}})\cdot{\mathbf{k}}_{\perp}\to\hbar\omega_{c}\begin{pmatrix}0&a^{\dagger}\\ a&0\end{pmatrix}$ (3)
[…]
Below we calculate the electric dipole moment of finite thickness films by endowing the LL eigenfunctions of the quasi-2D model Eq. (S1) with $z$-dependence by taking $C^{{\cal N},r}_{(i,\mu)}\rightarrow C^{{\cal N},r}_{(i,\mu)}\delta(z-z_{i})$, where $z_{i}\hat{{\mathbf{z}}}$ is the location of the $i$th surface with the center of the film taken as the origin. The macroscopic electronic polarization (dipole moment per volume) is:
[…]
where $\Omega=A(z_{2N-1}-z_{0})$ is the volume of the sample, $A$ is the film cross sectional area and we have noted that only negative energy states are occupied at neutrality. Here the $\Psi_{E}({\mathbf{r}})$ are real-space energy eigenfunctions that have support only within the sample. Taking account of the LL degeneracy $A/(2\pi l_{B}^{2})$ we obtain [34]
$P^{z}_{\text{el}}=-\frac{e^{2}B}{Ndhc}\sum_{i}\;z_{i}\;\Bigg{(}\sum_{E_{{\cal N},r}<0,\mu}|C^{{\cal N},r}_{(i,\mu)}|^{2}\Bigg{)}\equiv\alpha_{\text{me}}B^{z}.$ (5)
The factor in round brackets is the areal electronic charge density on surface $i$ and $d$ is the thickness of the unit cell for each layer. Since $m_{i}$ is odd under the mirror transformation $z\to-z$ [34], it yields a spin-polarization that is odd in $z$. In the ${\cal N}=0$ anomalous LL subspace, which contains only spin-up states, the mass terms give rise to a charge polarization. Because the block projected Hamiltonian is odd under the combined mirror and spin-reversal operations, the charge density contributed by the ${\cal N}\neq 0$ blocks is even under $z\to-z$ and there is no contribution to the dipole moment; the magnetoelectric response of our model comes entirely from the ${\cal N}=0$ block. Since the Hamiltonian of the ${\cal N}=0$ sector is a generalized SSH model with alternating hopping strengths $\Delta_{\text{S}}$ and $\Delta_{\text{D}}$, the quantized magnetoelectric coefficient $\alpha_{\text{me}}$ in our model follows from familiar analogous polarization properties of the SSH model. In the following, we detail how these properties depend on model parameters.
Bulk phase diagram and polarization properties— We can obtain the topological phase diagram of the bulk 3D Hamiltonian related to Eq. (S1) by tracking the band gap evolution when periodic boundary conditions are applied in $\hat{{\mathbf{z}}}$ direction  [^4] Technically, the model presented in Eq. (S1) is a quasi-2D low-energy effective ${\mathbf{k}}\cdot\bm{p}$ Hamiltonian. To generate a genuine 3D model for the electronic states in bulk, we need to replace the ${\mathbf{k}}\cdot\bm{p}$ basis states that the quasi-2D Hamiltonian is expressed in terms of with hybrid Wannier functions that are spatially localized in the ${\mathbf{z}}$ direction and the set of which contains one spin-up and one spin-down WF localized near each surface. We then apply periodic boundary conditions, by which we mean we take the model to describe an infinite chain of coupled-Dirac cone models (i.e., it involves a sum over all lattice vectors) in the ${\mathbf{z}}$-direction.. Our findings are shown in Fig. 2 (a). As explained below, these assignments have been confirmed by calculating the Zak phase of a corresponding one-dimensional (1D) model and also by explicitly evaluating the topological magnetoelectric coefficient of a 3D lattice regularization of Eq. (S1). The model parameters are chosen such that $\delta\equiv J_{\text{D}}/J_{\text{S}}\rightarrow 1$ corresponds to the non-magnetic case, for which a topological phase transition occurs when $|\Delta_{\text{D}}|=\Delta_{\text{S}}$. In this case there is no magnetic-field dependent charge transfer between layers in the bulk as illustrated by the grey curve in Fig. 2 (b). When $\delta\neq 1$ the interior layers are magnetized and magnetic-field dependent charge transfer does occur as illustrated by the green curve in Fig. 2 (b). Note that in the AFM case the topological phase boundary depends on $J_{\text{S}}$, that inter-layer charge transfer occurs in both the trivial ($\mathbb{Z}_{2}$-even) and non-trivial ($\mathbb{Z}_{2}$-odd) phases with magnitude dependent on model parameters, and that the rate of charge transfer with field jumps at the phase transition points, as shown in Fig. 2 (c) and (d).
Figure 2: Topological phase diagram and bulk charge redistribution. (a) Phase diagram of bulk AFM topological insulators vs. $\Delta_{\text{D}}/\Delta_{\text{S}}$ and $\delta\equiv J_{\text{D}}/J_{\text{S}}$ at $\Delta_{\text{S}}/J_{\text{S}}=1$. (The phase boundaries for $J_{\text{S}}=0.8\Delta_{\text{S}}$ and $J_{\text{S}}=1.5\Delta_{\text{S}}$ are plotted as purple and yellow dashed lines.) (b) Schematic magnetic-field induced charge density redistribution for non-magnetic (grey curve) and AFM (green curve) TIs; (c) and (d) Wavefunction amplitude variation (defined in (b)) vs. $\Delta_{\text{D}}$ for various values of $\delta$ when $J_{\text{S}}=1.5\Delta_{\text{S}}$ and $J_{\text{S}}=0.8\Delta_{\text{S}}$, respectively.
The magnetic-field dependent charge transfer seems to suggest that the magnetoelectric responses of both $\mathbb{Z}_{2}$-even and $\mathbb{Z}_{2}$-odd phases are non-zero, but this is not the case. As we have described near Eq. (3) and detailed in the Supplementary Material, the low-energy Hamiltonian in the 3D bulk can be block-diagonalized into subspaces that are uniquely identified by a LL index $\mathcal{N}$. As in finite samples, in this bulk model the magnetoelectric response comes entirely from the $\mathcal{N}=0$ LL subspace, within which the AFM Hamiltonian reduces to a family of identical 1D generalized SSH models – one for each state in the LL (see Supplementary Material):
$H_{\text{AFM}}(k_{z})$ $=(J_{\text{S}}-J_{\text{D}})\lambda_{z}\otimes\tau_{0}+\Delta_{\text{S}}\lambda_{0}\otimes\tau_{x}$
$+\Delta_{\text{D}}\lambda_{x}\otimes(\tau_{x}\cos k_{z}d+\tau_{y}\sin k_{z}d).$ (6)
In Eq. (6) each 1D unit cell has four sites since there are two van der Waals layers, each with top and bottom surface Dirac cones, and $\lambda_{i}$ and $\tau_{i}$ are Pauli matrices that act on the layer and surface degrees of freedom. (Spin degree of freedom doesn’t appear in Eq. (6) since states in the $\mathcal{N}=0$ LL subspace are spin-up polarized.) The ground state polarization in our low-energy 3D model has one bulk component, namely $P^{z}$, which can be calculated using expressions derived in the modern theory of polarization [38, 39]. The $\mathcal{N}=0$ contribution to $P^{z}$ is proportional to the 1D bulk polarization of Eq. (6) at half-filling: $P_{\text{SSH}}^{z}=e\sum_{n=1}^{2}\int_{-\pi/2d}^{\pi/2d}\frac{dk_{z}}{2\pi}\xi^{z}_{nn}(k_{z})$, where $\xi^{a}_{nm}(k_{z})\equiv i\langle u_{n,k_{z}}|\frac{\partial}{\partial k^{z}}u_{m,k_{z}}\rangle$ is the Berry connection in a (periodic) Hamiltonian gauge. $P_{\text{SSH}}^{z}$ is quantized because Eq. (6) has particle-hole symmetry: $U^{\dagger}_{\mathcal{C}}(k_{z})H_{\text{AFM}}(-k_{z})^{*}U_{\mathcal{C}}(k_{z})=-H_{\text{AFM}}(k_{z})$ for $U_{\mathcal{C}}(k_{z})=(\lambda_{x}\text{cos}dk_{z}-\lambda_{y}\text{sin}dk_{z})\otimes\tau_{z}$. Indeed, at half-filling this model can support topologically distinct ground states characterized by a $\mathbb{Z}_{2}$-valued invariant  [^5] See Sec. 2D of Ref. [6] and Ref. [58]., which is obtained by performing the 1D BZ integral in $P^{z}_{\text{SSH}}$. The association  [^6] See Sec. 2D of Ref. [6]. of $\mathbb{Z}_{2}$-even with $P_{\text{SSH}}^{z}=en$ and $\mathbb{Z}_{2}$-odd with $P_{\text{SSH}}^{z}=e(n+\frac{1}{2})$ for some $n\in\mathbb{Z}$ allows Fig. 2 (a) to be confirmed by a semi-analytic calculation (see Supplementary Material). A similar analysis applies in the non-magnetic case, for which the projected Hamiltonian in the $\mathcal{N}=0$ subspace is the usual SSH model. Thus $P^{z}$ does not distinguish between the microscopically distinct magnetic-field dependent charge redistribution (Fig. 2) in AFM and non-magnetic cases.
We can also evaluate the bulk magnetoelectric coefficient in Eq. (1) using a 3D tight-binding Hamiltonian [15] obtained by regularizing Eq. (S1) and employing the previously derived [6, 7] linear response expression. As described above, in band insulators with (generalized) TRS it has been argued that $\alpha^{il}=\delta^{il}\alpha_{\text{CS}}$, where $\alpha_{\text{CS}}$ is sensitive to the TRS-induced topological classification of the electronic ground state (encoded in a $\mathbb{Z}_{2}$ index). In general, $\alpha_{\text{CS}}$ is gauge-dependent and its explicit form employs a smooth global gauge of the vector bundle of occupied Bloch states over the 3D BZ torus. Thus, a smooth global Hamiltonian gauge $\mathfrak{u}$ identified pointwise by $\mathfrak{u}_{\bm{k}}=(\ket{u_{n\bm{k}}})_{n\in\{1,\ldots,N_{\text{occ}}\}}$ is viable (if it exists) and in that gauge
$\alpha^{\mathfrak{u}}_{\text{CS}}=\frac{-e^{2}}{2\hbar c}\epsilon^{abd}\int_{\text{BZ}}{\frac{d^{3}k}{(2\pi)^{3}}\left(\xi^{a}_{vv^{\prime}}\partial_{b}\xi^{d}_{v^{\prime}v}-\frac{2i}{3}\xi^{a}_{vv^{\prime}}\xi^{b}_{v^{\prime}v_{1}}\xi^{d}_{v_{1}v}\right)},$ (7)
where repeated indices are summed. (The implicit band sums run over the occupied bands $v,v^{\prime},v_{1}\in\{1,\ldots,N_{\text{occ}}\}$ only.) Although the expression for $\alpha^{\mathfrak{u}}_{\text{CS}}$ is gauge-dependent, transforming from one appropriate gauge  [^7] In this context one typically considers gauge transformations that decompose as a product of unitary transformations, one acting on the occupied states and another on the unoccupied states. A gauge choice that is periodic over BZ is always made. to another can only change its value by an integer multiple of $e^{2}/hc$ [6, 7].
Explicitly computing $\alpha_{\text{CS}}$ is challenging because a smooth global Hamiltonian gauge does not typically exist in $\mathbb{Z}_{2}$-odd insulators [43, 44], requiring a non-trivial Wannierization-like procedure to be employed. In our case, however, because the model has a generalized TRS and all energy bands appear in isolated degenerate pairs over the entire BZ, such a gauge does exist  [^8] In general, TRS implies that the (set of) Chern number(s) associated with each set of isolated bands vanishes, thus there exists a smooth global frame of the corresponding vector bundles. Due to the degeneracy in this special case, the components of each such frame are energy eigenvectors. (This same argument holds if there are no degeneracies between bands anywhere in BZ, but in the presence of a symmetry such as time-reversal or inversion, which ultimately gives rise to a $\mathbb{Z}_{2}$ index, this will never be the case.) Hence, a smooth global Hamiltonian gauge always exists in such circumstances.. In a related work [15] we explicitly evaluated Eq. (7) in the non-magnetic case. We can, at the cost of some additional complication, also find the analytic form of a smooth global Hamiltonian gauge in the AFM case. With this gauge in-hand, and employing the previously detailed [15] approximation that the basis of the tight-binding model (that results from lattice regularization) consists of atomic-like Wannier functions, we can semi-analytically evaluate Eq. (7). (See Supplementary Material.) This calculation again reproduces quantizated $\alpha_{\text{CS}}$ assignments in agreement with Fig. 2 (a).
Magnetoelectric response in thin films— In Fig. 3 (a) we show a $J_{\text{S}}=1.5\Delta_{\text{S}}$ phase diagram constructed by tracking thin film gap closings and confirmed by evaluating the magnetoelectric response coefficient $\alpha_{\text{me}}$, expressed here in units of $e^{2}/2hc$. The thin film results perfectly match the bulk $J_{\text{S}}=1.5\Delta_{\text{S}}$ results in Fig. 2. Notably, the thin film phase diagram has an extra boundary surrounding the region with $\alpha_{\text{me}}=2$, which results from the appearance of (top and bottom) surface Hall conductivities (Chern wrappers [14]) in this region of our model’s phase diagram, that are not related to bulk topology. The Chern wrapper region’s extent depends on details of the model that do not influence bulk topology; it is absent for $J_{\text{S}}\leq\Delta_{\text{S}}$, as illustrated in 3 (b). Note that the opportunity to identify bulk topological phase transitions by tracking thin film gaps is present only in magnetic $\mathbb{Z}_{2}$ TIs, which do not have gapless edge modes.
Figure 3: (a) Thin film phase diagram for $J_{\text{S}}=1.5\Delta_{\text{S}}$. The (quantized) magnetoelectric coefficient $\alpha_{\text{me}}$ is expressed here in units of $e^{2}/2hc$. The green region corresponds to $\mathbb{Z}_{2}$-odd and the red and grey regions to $\mathbb{Z}_{2}$-even bulk phases; the states in the red region have Chern wrappers as discussed in the main text, and therefore have $\alpha_{\text{me}}=2$. (b) Thin film phase diagram for $J_{\text{S}}=0.8\Delta_{\text{S}}$. (c) Schematic illustration of ${\cal N}=0$ (see main text) charge distributions (red curves) for $\alpha_{\text{me}}=0,\,1,\,2$ at $J_{\text{S}}=1.5\Delta_{\text{S}}$ and a small value of $\Delta_{\text{D}}$. In all cases there are both interior and surface contributions to $\alpha_{\text{me}}$ that are not quantized separately. The arrows indicate the orientation of the static magnetic moments in each van der Waals layer and the background color (white, blue, green) indicates the layer Chern number in the limit $\Delta_{\text{D}}\to 0$.
Figs. 3 (a) and (b) can be understood qualitatively by examining the $\Delta_{\text{D}}=0$ limit of the ${\cal N}=0$ Hamiltonian. In this hypothetical limit there is no hybridization between van der Waals layers and each layer is described by independent $2\times 2$ Hamiltonians with a surface degree-of-freedom:
$H_{\text{1L}}=\bar{m}\tau_{0}+\tilde{m}\tau_{z}+\Delta_{\text{S}}\tau_{x},$ (8)
where $\bar{m}\equiv(m_{t}+m_{b})/2$ and $\tilde{m}\equiv(m_{t}-m_{b})/2$ with $m_{t/b}$ the exchange coupling at top and bottom surface of each layer. On the outermost surfaces $m_{t/b}=\pm J_{\text{S}}$ whereas on the interior surfaces $m_{t/b}=\pm(1-\delta)J_{\text{S}}$, with the sign specified by the moment orientation in the same layer [34]. The eigenvalues of this Hamiltonian are $E=\bar{m}\pm(\Delta_{\text{S}}^{2}+\tilde{m}^{2})^{1/2}$. The phase diagram is controlled by the relative sizes of $|\bar{m}|$ and $(\Delta_{\text{S}}^{2}+\tilde{m}^{2})^{1/2}$ on the outermost and interior layers. In the $\alpha_{me}=0$ phase, $|\bar{m}|<(\Delta_{\text{S}}^{2}+\tilde{m}^{2})^{1/2}$ in all layers, in the $\alpha_{me}=1$ phase $|\bar{m}|>(\Delta_{\text{S}}^{2}+\tilde{m}^{2})^{1/2}$ in all layers, whereas in the $\alpha_{me}=2$ phase $|\bar{m}|$ is larger in the outside layers and smaller in the interior layers. In the $\alpha_{me}=1$ case, the ${\cal N}=0$ eigenfunctions occupy alternate layers as illustrated in the middle panel of Fig. 3 (c). For $\delta=0$ all layers are identical and
$\alpha_{\text{me}}=\sum_{j=1}^{N/2}(-1)^{N/2+j}\frac{2j-1}{N/2}=1$ (9)
is quantized even at finite $N$. More generally, for any values of $\Delta_{\text{D}}$ and $\delta$, the magnetoelectric coefficient achieves quantization only in sufficiently thick films [34].
Discussion— We have explored the magnetoelectric properties of layered $\mathbb{Z}_{2}$ topological insulators by evaluating the (electronic) polarization response of realistic thin film and bulk model Hamiltonians to a uniform dc magnetic field along the stacking axis. For an effective low-energy model we find that the polarization response arises entirely from the formation of $\mathcal{N}=0$ anomalous Landau levels described by 1D generalized SSH models. In the thin film case the total magnetoelectric response is quantized for sufficiently thick films, but it arises from non-zero and non-universal surface and interior contributions. (This behaviour is partially related to that of the alternating stacked Haldane model considered in Ref. [46].) We have reached these conclusions both by numerically examining the many-layer limit of the thin film calculations and by applying periodic boundary conditions in the stacking direction to the SSH models. Both approaches find that only the total polarization is quantized, in agreement with our 3D bulk calculations that relate the magnetoelectric response to the Chern-Simons 3-form [6, 7], which we evaluate explicitly for a lattice regularization of the low-energy model.
It is interesting to relate our results to efforts [47, 48] to express the the orbital magnetization ${\mathbf{M}}$ as a sum of contributions having either atomic-like ($\overline{{\mathbf{M}}}$) or itinerant ($\widetilde{{\mathbf{M}}}$) character. This partitioning can be made unambiguously, given a model with isolated energy bands that support localized Wannier functions, and associates $\overline{{\mathbf{M}}}$ with interior currents and $\widetilde{{\mathbf{M}}}$ with surface currents. However, when the bands are not isolated the partitioning is gauge-dependent. In previous work [15] on the non-magnetic case we were able, thanks to the analytic nature of our simplified model calculations, to demonstrate that there exists a smooth global (Hamiltonian) gauge $\mathfrak{u}$ in which the topological magnetoelectric response is entirely itinerant, i.e. $\widetilde{{\mathbf{M}}}_{\mathfrak{u}}^{(E)}=\alpha^{\mathfrak{u}}_{\text{CS}}{\mathbf{E}}$. We conjectured that this property is not specific to the simplified model, but that it is instead a bulk manifestation of the feature that in non-magnetic thin films magnetoelectric response can occur only near the surface, and as a bulk harbinger of the magnetic-dopant requirement for the topological magnetoelectric effect to manifest in non-magnetic TI films. In the Supplementary Material we explicitly demonstrate that in the AFM case there does not generically exist a smooth global gauge in which $\widetilde{{\mathbf{M}}}_{\mathfrak{u}}^{(E)}$ equals $\alpha^{\mathfrak{u}}_{\text{CS}}{\mathbf{E}}$. This difference between AFM and non-magnetic cases, which is encoded in the geometry  [^9] Recall that each global trivialization of a vector bundle corresponds with a global smooth frame (i.e., gauge) thereof (see, e.g., Chapter 10 of Ref. [LeeBook]). And although, for a given band insulator, the nonexistence of certain smooth global gauges has been related [44] to the topology of the vector bundle of occupied Bloch states $\mathcal{V}\rightarrow\text{BZ}$ [10], we do not suspect this to be the case here; we anticipate that in an AFM TI there could exist some smooth global gauges $\mathfrak{u}$ for which there exists a finely tuned point in the parameter space where ${\mathbf{M}}^{(E)}_{\mathfrak{u}}$ equals $\alpha_{\text{CS}}^{\mathfrak{u}}{\mathbf{E}}$. Thus, whether such a gauge generically exists is taken to be a geometric property of $\mathcal{V}$. of the corresponding vector bundles of occupied Bloch states, can be understood as a bulk manifestation of magnetoelectric response occurring both in the interior and at the surface of AFM TIs. Thus, although $\alpha_{\text{CS}}$ is determined by the (generalized) TRS-induced $\mathbb{Z}_{2}$ topological classification of the electronic ground state, which does not distinguish between non-magnetic and AFM TIs, the nature of that response – whether it is entirely itinerant or not – does distinguish them.
Our results provide a conditional answer to the thorny conceptual issue as to whether the topological magnetoelectric effect should be viewed as an interior or surface effect. We have previously emphasized that the magnetoelectric response of ordinary non-magnetic TIs is localized at their surfaces, and is realized only when the surfaces are gapped by local magnetic dopants. When the bulk is magnetic, surface dopants are no longer required and a response in the interior is allowed by symmetry. Our finding that the interior response does occur but is not (separately) quantized, shows that it is the total response that is related to bulk topology, and that the mechanism by which it is manifested generically involves both interior and surface contributions. Indeed, we expect that this periodic microscopic scale magnetic-field-induced charge density variation that appears within AFM TIs (and is absent in non-magnetic TIs) can be observed using high-sensitivity high-resolution X-ray imaging techniques [50, 51, 52, 53, 54, 55], and that these measurements might be aided by the fact that the density varies only in the magnetic field direction. The inability to assign a quantized response of a topological state wholly to interior or surface is analogous to the related issue of non-universal interior-edge current partition [16, 17, 18, 19] in the IQHE. In both cases, bulk topology constrains the total response but not its partitioning.
Acknowledgements. This work was supported by Welch Foundation grant F-2112.
Supplementary Material for ’Non-universal surface magnetoelectric response in antiferromagnetic topological insulators”
Appendix S1 Landau level representation
A previously developed [30] coupled-Dirac cone model is used to describe the low-energy electronic states in the layered compounds that are of interest in this paper. This model retains two Dirac cones in each layer as low-energy electronic degrees of freedom and the total effective Hamiltonian is
$\hat{H}=\sum_{{\mathbf{k}}_{\perp},ij,\mu\gamma}\Big{[}\Big{(}\,(-1)^{i}\hbar v_{D}(\hat{z}\times{\mathbf{\sigma}})\cdot{\mathbf{k}}_{\perp}+m_{i}\sigma_{z}\Big{)}_{\mu,\gamma}\delta_{ij}+\Delta_{ij}(1-\delta_{ij})\delta_{\mu,\gamma}\Big{]}\hat{c}_{{\mathbf{k}}_{\perp},i,\mu}^{\dagger}\hat{c}_{{\mathbf{k}}_{\perp},j,\gamma}.$ (S1)
Here ${\mathbf{k}}_{\perp}\equiv(k_{x},k_{y},0)$ for stacking axis $\hat{\bm{z}}$, $i,j\in\{0,\ldots,2N-1\}$ where $i=2(l-1)$ $(i=2l-1)$ label states that are associated with the bottom (top) surface of layer $l\in\{1,\ldots,N\}$ and $\mu,\gamma\in\{\uparrow,\downarrow\}$ label up and down spin components of $\sigma_{z}$, $\Delta_{ij}$ are intra- and inter-layer hybridization parameters, $v_{{}_{D}}$ is the velocity of the isolated Dirac cones, and $m_{i}$ are mass terms that account for exchange coupling of the Dirac-like states at surface $i$ to static local magnetic moments (taken aligned with the stacking direction) that can be present in each layer. We set $m_{i}=\sum_{l}J_{il}M_{l}$ with dimensionless magnetization $M_{l}=0$ if layer $l$ is non-magnetic layers and ±$\pm 1$ if layer $l$ is magnetic. As mentioned in the main text, we employ the usual prescription to account for an external magnetic field ${\mathbf{B}}=-B\hat{{\mathbf{z}}}$ in a ${\mathbf{k}}\cdot\bm{p}$ model: $\hbar{\mathbf{k}}_{\perp}\rightarrow-i\hbar(\partial/\partial x,\,\partial/\partial y,0)$ and choose the Landau gauge ${\mathbf{A}}=(0,-Bx,0)$. Under this recipe the Dirac Hamiltonian for each surface transforms as $H_{D}({\mathbf{k}}_{\perp})=\hbar v_{D}(\sigma_{x}k_{y}-\sigma_{y}k_{x})\rightarrow H_{D}(x,y)\equiv\hbar v_{D}\big{(}-\sigma_{x}(i\frac{\partial}{\partial y}+x/l_{B}^{2})+i\sigma_{y}\frac{\partial}{\partial x}\big{)}$ with $l_{B}\equiv\sqrt{\hbar c/eB}$. The total Hamiltonian that is obtained from employing this procedure in Eq. (S1) has translational symmetry in $\hat{{\mathbf{y}}}$, thus energy eigenfunctions can be written as
$\Psi_{E,q_{y}}(x,y)=\frac{e^{iq_{y}y}}{\sqrt{L_{y}}}\Phi_{E,q_{y}}(x)\equiv\frac{e^{iq_{y}y}}{\sqrt{L_{y}}}\prod_{i=0}^{2N-1}\Big{(}\phi_{E,q_{y},(i,\uparrow)}(x),\,\phi_{E,q_{y},(i,\downarrow)}(x)\Big{)}.$ (S2)
Since the only position dependence of the total Hamiltonian appears via the terms resembling $H_{D}$ for each surface, we focus on those contributions first. For an isolated surface of a non-magnetic layer the energy eigenfunctions satisfy $H_{D}^{2}(x,y)\Psi_{E,q_{y}}(x,y)=E^{2}\Psi_{E,q_{y}}(x,y)$ and $\Phi_{E,q_{y}}(x)=\big{(}\phi_{E,q_{y},\uparrow}(x),\,\phi_{E,q_{y},\downarrow}(x)\big{)}$, where
$(\hbar v_{D})^{2}(\partial_{x}+q_{y}-x/l_{B}^{2})(-\partial_{x}+q_{y}-x/l_{B}^{2})\phi_{E,q_{y},\uparrow}(x)$ $=E^{2}\phi_{E,q_{y},\uparrow}(x),$
$(\hbar v_{D})^{2}(-\partial_{x}+q_{y}-x/l_{B}^{2})(\partial_{x}+q_{y}-x/l_{B}^{2})\phi_{E,q_{y},\downarrow}(x)$ $=E^{2}\phi_{E,q_{y},\downarrow}(x).$ (S3)
With the substitutions $\tilde{x}\equiv l_{B}q_{y}-x/l_{B}$ and $\partial_{\tilde{x}}=-l_{B}\partial_{x}$, define the usual harmonic oscillator creation and annihilation differential operators $a^{\dagger}\equiv(\tilde{x}-\partial_{\tilde{x}})/\sqrt{2}$ and $a\equiv(\tilde{x}+\partial_{\tilde{x}})/\sqrt{2}$, which satisfy $[a,a^{\dagger}]=1$. Then, Eq. (S3) is recast as
$2\left(\frac{\hbar v_{D}}{l_{B}}\right)^{2}\begin{pmatrix}a^{\dagger}a&0\\ 0&aa^{\dagger}\end{pmatrix}\Phi_{E,q_{y}}(x)\equiv H^{2}_{D}(x,\partial_{x};q_{y})\Phi_{E,q_{y}}(x)=E^{2}\Phi_{E,q_{y}}(x).$ (S4)
From the usual analysis of the harmonic oscillator  [^10] See, e.g., pg. 89-94 of Ref. [57]. we have that eigenfunctions of $a^{\dagger}a$ are
$\chi_{n}(\tilde{x})=H_{n}(x-q_{y}l_{B}^{2})e^{-(x-q_{y}l_{B}^{2})^{2}/4l_{B}^{2}},$ (S5)
where $H_{n}(x)$ is the Hermite polynomial $H_{n}(x)=(-1)^{n}e^{x^{2}}\frac{d^{n}}{dx^{n}}e^{-x^{2}}$, and have corresponding eigenvalues $n$. These functions satisfy $a^{\dagger}\chi_{n}(\tilde{x})\propto\chi_{n+1}(\tilde{x})$ and $a\chi_{n+1}(\tilde{x})\propto\chi_{n}(\tilde{x})$ for integers $n\geq 0$. Using this we find that solutions of the eigenfunction-eigenvalue equation (S4) can be written in the general form
$\Phi_{E,q_{y}}(x)\equiv\Phi_{n}(\tilde{x})=\begin{cases}\big{(}C^{n}_{\uparrow}\chi_{n}(\tilde{x}),\,C^{n}_{\downarrow}\chi_{n-1}(\tilde{x})\big{)},&\text{ for }n\neq 0\\ \big{(}C^{0}_{\uparrow}\chi_{0}(\tilde{x}),\,0\big{)},&\text{ for }n=0\end{cases}$ (S6)
for $C^{n}_{\mu}\in\mathbb{C}$ and the corresponding eigenvalues are $E_{n}^{2}=2n(\hbar v_{D}/l_{B})^{2}$. This implies $E_{n}=\pm\hbar\omega_{c}\sqrt{n}$ where $\omega_{c}\equiv\sqrt{2}v_{D}/l_{B}$. Indeed, these general forms can diagonalize the isolated non-magnetic surface Dirac Hamiltonian
$H_{D}({\mathbf{k}}_{\perp})\rightarrow H_{D}(x,\partial_{x};q_{y})\equiv\hbar\omega_{c}\begin{pmatrix}0&a^{\dagger}\\ a&0\end{pmatrix}.$ (S7)
In the total Hamiltonian (for a $N$ layer film) there is a contribution related to $H_{D}$ for each surface $i$ that couples $(\chi_{n}(\tilde{x}),0)$ only to $(0,\chi_{n-1}(\tilde{x}))$. A reasonable ansatz for an energy eigenfunction could therefore involve $\Phi_{E}(\tilde{x})$ of the form $\prod_{i=0}^{2N-1}\big{(}C^{n_{i}}_{(i,\uparrow)}\chi_{n_{i}}(\tilde{x}),\,C^{n_{i}}_{(i,\downarrow)}\chi_{n_{i}-1}(\tilde{x})\big{)}$, where $n_{i}\geq 0$ for all surfaces $i\in\{1,\ldots,2N-1\}$. However, since there are no other terms in (S1) that involve ${\mathbf{k}}_{\perp}$, there are no other terms in the Hamiltonian (after $\hbar{\mathbf{k}}_{\perp}\rightarrow-i\hbar(\partial/\partial x,\,\partial/\partial y)$) that involve $a$ or $a^{\dagger}$. That is, when acting on the ansatz eigenfunction the other terms in the Hamiltonian can map a component at surface $i$ into a component at surface $j$, but only in such a way that the new component at surface $j$ is in the space spanned by $\chi_{n_{i}}(\tilde{x})$. In our model the states at each surface are coupled to those at two surfaces adjacent to it (via $\Delta_{\text{S}}$ and $\Delta_{\text{D}}$), thus the ansatz is valid only if $n_{1}=\ldots=n_{2N-1}\equiv n$ for $n\geq 0$. (We sometimes adopt the notation $\mathcal{N}$ for this $n$.) The total energy eigenfunctions can therefore be written in the general form
$\Psi_{n,q_{y}}(x,y)=\begin{cases}\frac{e^{iq_{y}y}}{\sqrt{L_{y}}}\prod_{i=0}^{2N-1}\big{(}C^{n}_{(i,\uparrow)}\chi_{n}(\tilde{x}),\,C^{n}_{(i,\downarrow)}\chi_{n-1}(\tilde{x})\big{)},&\text{ for }n\neq 0\\ \frac{e^{iq_{y}y}}{\sqrt{L_{y}}}\prod_{i=0}^{2N-1}\big{(}C^{0}_{(i,\uparrow)}\chi_{0}(\tilde{x}),\,0\big{)},&\text{ for }n=0\end{cases}.$ (S8)
S1.1 Multi-layer antiferromagnetic materials
In this section we specify the action of the Hamiltonian (for a material consisting of $N$ layers) in the subspace defined by LL index $n$ via a matrix acting on a collection of eigenfunction components $C^{n}_{(i,\mu)}$ introduced in Eq. (S8). For $n\neq 0$ the Hamiltonian matrix acts on the collection of components $\prod_{i=0}^{2N-1}\big{(}C^{n}_{(i,\uparrow)},\,C^{n}_{(i,\downarrow)}\big{)}$. The bulk ($N\rightarrow\infty$) multi-layer antiferromagnetic Hamiltonian matrix for $n\neq 0$ is
$H_{\text{bulk}}^{n}=\left(\begin{array}[]{cccccccc}J_{\text{S}}-J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0&0&0&\Delta_{\text{D}}e^{-2ik_{z}d}&0\\ \sqrt{n}\hbar\omega_{c}&-J_{\text{S}}+J_{\text{D}}&0&\Delta_{\text{S}}&0&0&0&\Delta_{\text{D}}e^{-2ik_{z}d}\\ \Delta_{\text{S}}&0&J_{\text{S}}-J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{D}}&0&0&0\\ 0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&-J_{\text{S}}+J_{\text{D}}&0&\Delta_{\text{D}}&0&0\\ 0&0&\Delta_{\text{D}}&0&-J_{\text{S}}+J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0\\ 0&0&0&\Delta_{\text{D}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}-J_{\text{D}}&0&\Delta_{\text{S}}\\ \Delta_{\text{D}}e^{2ik_{z}d}&0&0&0&\Delta_{\text{S}}&0&-J_{\text{S}}+J_{\text{D}}&\sqrt{n}\hbar\omega_{c}\\ 0&\Delta_{\text{D}}e^{2ik_{z}d}&0&&0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}-J_{\text{D}}\\ \end{array}\right).$ (S9)
For a thin film, for example, a film consisting of two layers ($N=2$), the Hamiltonian matrix is
$H_{\text{2L}}^{n}=\left(\begin{array}[]{cccccccc}J_{\text{S}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0&0&0&0&0\\ \sqrt{n}\hbar\omega_{c}&-J_{\text{S}}&0&\Delta_{\text{S}}&0&0&0&0\\ \Delta_{\text{S}}&0&J_{\text{S}}-J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{D}}&0&0&0\\ 0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&-J_{\text{S}}+J_{\text{D}}&0&\Delta_{\text{D}}&0&0\\ 0&0&\Delta_{\text{D}}&0&-J_{\text{S}}+J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0\\ 0&0&0&\Delta_{\text{D}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}-J_{\text{D}}&0&\Delta_{\text{S}}\\ 0&0&0&0&\Delta_{\text{S}}&0&-J_{\text{S}}&\sqrt{n}\hbar\omega_{c}\\ 0&0&0&&0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}\\ \end{array}\right).$ (S10)
Under the mirror transformation $z\rightarrow-z$, which we encode in a unitary operator $M$ that maps the spin-$\sigma$ orbital at the surface located at $z_{i}\hat{{\mathbf{z}}}$ to the spin-$\sigma$ orbital located at $-z_{i}\hat{{\mathbf{z}}}$ (thus, $M=M^{-1}$), we find
$MH_{\text{2L}}^{n}M=\left(\begin{array}[]{cccccccc}-J_{\text{S}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0&0&0&0&0\\ \sqrt{n}\hbar\omega_{c}&J_{\text{S}}&0&\Delta_{\text{S}}&0&0&0&0\\ \Delta_{\text{S}}&0&-J_{\text{S}}+J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{D}}&0&0&0\\ 0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}-J_{\text{D}}&0&\Delta_{\text{D}}&0&0\\ 0&0&\Delta_{\text{D}}&0&J_{\text{S}}-J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0\\ 0&0&0&\Delta_{\text{D}}&\sqrt{n}\hbar\omega_{c}&-J_{\text{S}}+J_{\text{D}}&0&\Delta_{\text{S}}\\ 0&0&0&0&\Delta_{\text{S}}&0&J_{\text{S}}&\sqrt{n}\hbar\omega_{c}\\ 0&0&0&&0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&-J_{\text{S}}\\ \end{array}\right).$ (S11)
When composed with a spin-flip operation each $n\neq 0$ Hamiltonian is invariant, thus the spin-flip times mirror operator is a symmetry.
We can similarly consider the multi-layer antiferromagnetic Hamiltonian for $n=0$, in which case the matrix acts on the collection of components $\prod_{i=0}^{2N-1}\big{(}C^{n}_{(i,\uparrow)}\big{)}$. In the bulk case, the Hamiltonian matrix is
$H^{n=0}_{\text{bulk}}=\left(\begin{array}[]{cccc}J_{\text{S}}-J_{\text{D}}&\Delta_{\text{S}}&0&\Delta_{\text{D}}e^{-2ik_{z}d}\\ \Delta_{\text{S}}&J_{\text{S}}-J_{\text{D}}&\Delta_{\text{D}}&0\\ 0&\Delta_{\text{D}}&-J_{\text{S}}+J_{\text{D}}&\Delta_{\text{S}}\\ \Delta_{\text{D}}e^{2ik_{z}d}&0&\Delta_{\text{S}}&-J_{\text{S}}+J_{\text{D}}\\ \end{array}\right).$ (S12)
In the case of a 2-layer thin film the Hamiltonian matrix is
$H^{n=0}_{\text{2L}}=\left(\begin{array}[]{cccc}J_{\text{S}}&\Delta_{\text{S}}&0&0\\ \Delta_{\text{S}}&J_{\text{S}}-J_{\text{D}}&\Delta_{\text{D}}&0\\ 0&\Delta_{\text{D}}&-J_{\text{S}}+J_{\text{D}}&\Delta_{\text{S}}\\ 0&0&\Delta_{\text{S}}&-J_{\text{S}}\\ \end{array}\right).$ (S13)
In contrast to the $n\neq 0$ case, $H^{n=0}_{\text{2L}}$ is not symmetric under the spin-flip times mirror transformation.
At $k_{z}=0$ the energy eigenvalues of $H^{n=0}_{\text{bulk}}$ (Eq. (S12)) are:
$E_{1}$ $=-E_{4}=\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}+\Delta_{\text{S}},$
$E_{2}$ $=-E_{3}=\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}-\Delta_{\text{S}},$ (S14)
with corresponding (non-normalized) eigenvectors:
$\Psi_{1}$ $=\left(\frac{J_{\text{S}}-J_{\text{D}}+\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}}{\Delta_{\text{D}}},\frac{J_{\text{S}}-J_{\text{D}}+\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}}{\Delta_{\text{D}}},1,1\right),$
$\Psi_{2}$ $=\left(\frac{J_{\text{S}}-J_{\text{D}}+\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}}{\Delta_{\text{D}}},-\frac{J_{\text{S}}-J_{\text{D}}+\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}}{\Delta_{\text{D}}},-1,1\right),$
$\Psi_{3}$ $=\left(\frac{J_{\text{S}}-J_{\text{D}}-\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}}{\Delta_{\text{D}}},\frac{J_{\text{S}}-J_{\text{D}}-\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}}{\Delta_{\text{D}}},1,1\right),$
$\Psi_{4}$ $=\left(\frac{J_{\text{S}}-J_{\text{D}}-\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}}{\Delta_{\text{D}}},-\frac{J_{\text{S}}-J_{\text{D}}-\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}}{\Delta_{\text{D}}},-1,1\right).$ (S15)
At half-filling the bulk band gap closes when $E_{2}=E_{3}=0$, which occurs only if $\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}=\Delta_{\text{S}}$. Around this phase transition point, the eigenstate $\Psi_{2}$ or $\Psi_{3}$ is occupied, leading to a discontinuous change of the amplitude of the wavefunctions for occupied states.
S1.2 Bulk polarization of the electronic ground state in the low-energy model
In the main text we demonstrated that, due to the symmetry of Eq. (S9), the $\mathcal{N}>0$ Landau level subspaces do not contribute to the electronic polarization. Thus we restrict focus to the Hamiltonian in the $\mathcal{N}=0$ subspace, Eq. (S12), which takes the form of a generalized SSH model. That is, when we consider an infinite number of layers (i.e. implement periodic boundary conditions) in Eq. (S1), the result takes a tight-binding form [15] in the stacking ($\hat{\bm{z}}$) direction. And when this 3D effective model is projected into the $\mathcal{N}=0$ subspace, the result is a one-dimensional (1D) tight-binding model. The ground state polarization in our low-energy 3D model has one bulk component, namely $P^{z}$, which can be calculated using expressions derived in the modern theory of polarization [38, 39]. The $\mathcal{N}=0$ contribution to $P^{z}$ is proportional to the 1D bulk polarization of Eq. (S12) at half-filling, namely
$P_{\text{SSH}}^{z}=e\sum_{n=1}^{2}\int_{\text{BZ}}\frac{dk_{z}}{2\pi}\xi^{z}_{nn}(k_{z}),$ (S16)
where $\xi^{z}_{nm}(k_{z})\equiv i\innerproduct{u_{n,k_{z}}}{\frac{\partial}{\partial k^{z}}u_{m,k_{z}}}$ are band components of the Berry connection induced by a choice of Bloch energy eigenvectors with cell-periodic parts |$\ket{u_{n,k_{z}}}$. Inversion symmetry of the model dictates that $P_{\text{SSH}}^{z}\text{ mod }e\in\{0,e/2\}$ [60]. A set of analytical eigenvectors of Eq. (S12) can be found and the components differentiated with respect to $k_{z}$ in an effort to obtain $\xi^{z}_{nn}$.
One issue remains: how do we take the $k_{z}$-derivative of the basis states with respect to which the Hamiltonian matrix (S12) is written? Typically in phenomenological tight-binding models those basis states are taken to be such that the corresponding orbitals are highly localized in real-space and centered at the ionic positions. Often the derivatives of these basis states do not affect the results of calculations of topological quantities (e.g., polarization of the usual SSH model or the Chern number of the Haldane model) and one can get away with taking the basis states to be independent of $\bm{k}$. However, in the present case accounting for these derivatives becomes important. For if we assume the basis states to be independent of $k_{z}$ then we find, in general, that $P_{\text{SSH}}^{z}\text{ mod }e\notin\{0,e/2\}$. Thus, that assumption leads to a contradiction and is therefore not valid. Instead, we employ approximation Eq. (S29) (detailed below) and take $\delta^{z}_{0}=-\delta^{z}_{3}=-d$ and $\delta^{z}_{1}=-\delta^{z}_{2}=0$; we assign the spin-up orbital associated with the bottom (top) surface of the bottom (top) layer the position $z=-d$ ($z=d$) and the spin-up orbital associated with the top (bottom) surface of the bottom (top) layer the position $z=0$ in the unit cell whose center is taken to be the point in the middle of the two layers (which are stacked on top of one another, see Fig. 1 of Ref. [30]). With this, we semi-analytically compute (S16), the results of which are in perfect agreement with Fig. 2 (a) of the main text. It is notable that in this case the calculation of $P^{z}$ is sensitive to the approximations made to the basis functions, while later in calculations of the topological magnetoelectric coefficient, this is no longer the case.
S1.3 Antiferromagnetic thin films
The phase diagram in Fig. 3 of the main text may be understood in the limit of $\Delta_{\text{D}}\rightarrow 0$. In this case there is no hybridization between electronic states in different van der Waals layers and within each layer the $\mathcal{N}=0$ Hamiltonian generally acts on the surface degree of freedom via a $2\times 2$ matrix:
$H_{\text{1L}}^{n=0}=\left(\begin{array}[]{cccc}m_{t}&\Delta_{\text{S}}\\ \Delta_{\text{S}}&m_{b}\\ \end{array}\right)=\frac{m_{t}+m_{b}}{2}\left(\begin{array}[]{cccc}1&0\\ 0&1\\ \end{array}\right)+\left(\begin{array}[]{cccc}(m_{t}-m_{b})/2&\Delta_{\text{S}}\\ \Delta_{\text{S}}&-(m_{t}-m_{b})/2\\ \end{array}\right),$ (S17)
which may be written as $H_{\text{1L}}=\bar{m}\tau_{0}+\tilde{m}\tau_{z}+\Delta_{\text{S}}\tau_{x}$ if we define $\bar{m}\equiv(m_{t}+m_{b})/2$ and $\tilde{m}\equiv(m_{t}-m_{b})/2$. Here $\tau_{i}$ denote the Pauli matrices. On the outside surfaces $m_{t/b}=\pm J_{\text{S}}$ whereas on the interior surfaces $m_{t/b}=\pm(1-\delta)J_{\text{S}}$ with $\delta\equiv J_{\text{D}}/J_{\text{S}}$. In the limit of $\delta\rightarrow 0$, the Hamiltonian in the $\mathcal{N}=0$ subspace have for a two layer thin film is:
$H_{\text{2L}}^{n=0}=\left(\begin{array}[]{cccc}J_{\text{S}}&\Delta_{\text{S}}&0&0\\ \Delta_{\text{S}}&J_{\text{S}}&0&0\\ 0&0&-J_{\text{S}}&\Delta_{\text{S}}\\ 0&0&\Delta_{\text{S}}&-J_{\text{S}}\\ \end{array}\right).$ (S18)
The eigenvalues are ±$\pm J_{\text{S}}\pm\Delta_{\text{S}}$ and thus have a phase transition at $J_{\text{S}}=\Delta_{\text{S}}$. If $J_{\text{S}}<\Delta_{\text{S}}$, the two states with negative energies are $E_{-}=\pm J_{\text{S}}-\Delta_{\text{S}}$ which belong different layers, else if $J_{\text{S}}>\Delta_{\text{S}}$, the two states with negative energies are $E_{-}=-J_{\text{S}}\pm\Delta_{\text{S}}$ which belong to the same layer.
In the case of $\delta\neq 0$, the layers in $N$-layer thin films may be grouped into the outer layers and interior layers. As an illustration, the following Hamiltonian describes a 4-layer thin film:
$H^{n=0}_{\text{4L}}=\left(\begin{array}[]{cccccccc}J_{\text{S}}&\Delta_{\text{S}}&0&0&0&0&0&0\\ \Delta_{\text{S}}&J_{\text{S}}-J_{\text{D}}&0&0&0&0&0&0\\ 0&0&-J_{\text{S}}+J_{\text{D}}&\Delta_{\text{S}}&0&0&0&0\\ 0&0&\Delta_{\text{S}}&-J_{\text{S}}+J_{\text{D}}&0&0&0&0\\ 0&0&0&0&J_{\text{S}}-J_{\text{D}}&\Delta_{\text{S}}&0&0\\ 0&0&0&0&\Delta_{\text{S}}&J_{\text{S}}-J_{\text{D}}&0&0\\ 0&0&0&0&0&0&-J_{\text{S}}+J_{\text{D}}&\Delta_{\text{S}}\\ 0&0&0&0&0&0&\Delta_{\text{S}}&-J_{\text{S}}\\ \end{array}\right).$ (S19)
Generalized to $N$-layer thin films, the energy eigenvalues and (non-normalized) eigenstates can be solved as:
$E_{1}=(1-\delta/2)J_{\text{S}}+\sqrt{\Delta_{\text{S}}^{2}+(J_{\text{S}}\delta/2)^{2}};$ $\Psi_{1}=\big{(}\phi_{+},1,0,0,\cdots,0,0,0,0\big{)},$
$E_{2}=(1-\delta/2)J_{\text{S}}-\sqrt{\Delta_{\text{S}}^{2}+(J_{\text{S}}\delta/2)^{2}};$ $\Psi_{2}=\big{(}\phi_{-},1,0,0,\cdots,0,0,0,0\big{)},$
$E_{3}=-(1-\delta)J_{\text{S}}+\Delta_{\text{S}};$ $\Psi_{3}=\big{(}0,0,1,1,\cdots,0,0,0,0\big{)},$
$E_{4}=-(1-\delta)J_{\text{S}}-\Delta_{\text{S}};$ $\Psi_{4}=\big{(}0,0,-1,1,\cdots,0,0,0,0\big{)},$
[…]
$E_{2N-3}=(1-\delta)J_{\text{S}}+\Delta_{\text{S}};$ $\Psi_{2N-3}=\big{(}0,0,0,0,\cdots,1,1,0,0\big{)},$
$E_{2N-2}=(1-\delta)J_{\text{S}}-\Delta_{\text{S}};$ $\Psi_{2N-2}=\big{(}0,0,0,0,\cdots,-1,1,0,0\big{)},$
$E_{2N-1}=-(1-\delta/2)J_{\text{S}}+\sqrt{\Delta_{\text{S}}^{2}+(J_{\text{S}}\delta/2)^{2}};$ $\Psi_{2N-1}=\big{(}0,0,0,0,\cdots,0,0,\phi_{+},1\big{)},$
$E_{2N}=-(1-\delta/2)J_{\text{S}}-\sqrt{\Delta_{\text{S}}^{2}+(J_{\text{S}}\delta/2)^{2}};$ $\Psi_{2N}=\big{(}0,0,0,0,\cdots,0,0,\phi_{-},1\big{)}.$ (S20)
Here $\phi_{\pm}\equiv J_{\text{S}}\delta/2\Delta_{\text{S}}\pm\sqrt{1+(J_{\text{S}}\delta/2\Delta_{\text{S}})^{2}}$, the normalization factor of the wavefunctions are $\mathcal{A}_{\pm}=1/\sqrt{1+|\phi_{\pm}|^{2}}$ and $1/\sqrt{2}$. Note that in the limit of $\delta\to 0$, $\phi_{\pm}\to\pm 1$ and $\mathcal{A}_{\pm}\to 1/\sqrt{2}$. When setting the center of the thin films as the origin ($z=0$) and ascribing to the top and bottom surface of each layer the position $z_{j}=jd_{s}+(j-1/2)d_{v}$ and $z_{j}=(j-1)d_{s}+(j-1/2)d_{v}$ for $z>0$, whereas $z_{j}=-jd_{s}-(j-1/2)d_{v}$ and $z_{j}=-(j-1)d_{s}-(j-1/2)d_{v}$ for $z<0$, with $d_{s}$ the thickness of each layer (distance between the top and bottom surface for each layer) and $d_{v}$ the distance between each layer. Here $j=1,2,\ldots,N/2$ where $N$ is the total number of layers of the thin films.
As shown in the main text, the magnetoelectric coefficient may be written as:
$\alpha_{\text{me}}=\frac{e^{2}}{N(d_{s}+d_{v})hc}\sum_{j,E<E_{F}}|C_{E}(z_{j})|^{2}z_{j}.$ (S21)
Here $E_{F}$ is the Fermi energy which is 00 in our model. In the following we will discuss the magnetoelectric coefficient for parameters in different regions:
  1. 1.
    $J_{\text{S}}<\Delta_{\text{S}}$
    For $J_{\text{S}}<\Delta_{\text{S}}$, $E_{2n}<0$ in Eq. (S1.3) and thus the contribution of interior layers is 0 since $|C_{E}(z_{j})|^{2}=C_{E}(-z_{j})|^{2}$.
    $\alpha_{\text{me}}|_{J_{\text{S}}<\Delta_{\text{S}}}=\frac{e^{2}}{Nhc}\frac{d_{s}}{d_{s}+d_{v}}\mathcal{A}_{-}^{2}(|\phi_{-}|^{2}-1),$ (S22)
    which is always equal to 00 when $N\to\infty$ or $\delta\to 0$.
    Figure S1: (a) Thin film phase diagram for $J_{\text{S}}=1.5\Delta_{\text{S}}$ as discussed in the main text. (b)-(d) are the plots of the magnetoelectric coefficient versus the thickness of thin films for various $\delta$ and $\Delta_{\text{D}}$, with the values of $\delta$ labeled with black dashed lines in (a).
  2. 2.
    $J_{\text{S}}>\Delta_{\text{S}}$
    For the case of $J_{\text{S}}>\Delta_{\text{S}}$, as we show in the main text there are two phase transition points when varying the value of $\delta$, it is easy to see in the band gap which closes two times.
    1. (a)
      $\delta<1-\Delta_{\text{S}}/J_{\text{S}}$
      In this case $(1-\delta/2)J_{\text{S}}-\sqrt{\Delta_{\text{S}}^{2}+(J_{\text{S}}\delta/2)^{2}}>0$ and $(1-\delta)J_{\text{S}}-\Delta_{\text{S}}>0$, layers with even number is occupied, and thus the contribution from the interior layers is $\alpha_{\text{me}}^{j}=e^{2}(2j-1)/Nhc$. For outer layers, the contribution is:
      $\alpha_{\text{me}}^{\text{out}}=2\frac{e^{2}}{hc}\frac{N-1}{2N}+\frac{d_{s}}{2N(d_{s}+d_{v})}(\mathcal{A}_{+}^{2}+\mathcal{A}_{-}^{2}-\mathcal{A}_{+}^{2}\phi_{+}^{2}-\mathcal{A}_{-}^{2}\phi_{-}^{2}),$ (S23)
      where we used the fact that $\mathcal{A}_{+}^{2}+\mathcal{A}_{-}^{2}+\mathcal{A}_{+}^{2}\phi_{+}^{2}+\mathcal{A}_{-}^{2}\phi_{-}^{2}=2$ since $\mathcal{A}_{\pm}=1/\sqrt{1+\phi_{\pm}^{2}}$. The total magnetoelectric coefficient is thus:
      $\alpha_{\text{me}}|_{J_{\text{S}}>\Delta_{\text{S}},\delta<1-\Delta_{\text{S}}/J_{\text{S}}}=\alpha_{\text{me}}^{\text{out}}+\sum_{j=1}^{N/2-1}(-1)^{N/2+j}\alpha_{\text{me}}^{j}.$ (S24)
      Here the factor $(-1)^{N/2+j}$ is introduced so that the sign of $\alpha_{\text{me}}$ keeps the same different thickness of thin films. In the limit of $\delta\to 0$, the magnetoelectric coefficient becomes:
      $\alpha_{\text{me}}|_{J_{\text{S}}>\Delta_{\text{S}},\delta\to 0}=\frac{e^{2}}{2hc}\sum_{j=1}^{N/2}(-1)^{N/2+j}\frac{2j-1}{N/2}.$ (S25)
    2. (b)
      $1-\Delta_{\text{S}}/J_{\text{S}}<\delta<1-(\Delta_{\text{S}}/J_{\text{S}})^{2}$
      In this case, $(1-\delta/2)J_{\text{S}}-\sqrt{\Delta_{\text{S}}^{2}+(J_{\text{S}}\delta/2)^{2}}>0$ and $(1-\delta)J_{\text{S}}-\Delta_{\text{S}}<0$, the interior layer contribution is zero as $|C_{E}(z_{j})|^{2}=C_{E}(-z_{j})|^{2}$. The contribution to the magnetoelectric coefficient is thus only from the outer layer:
      $\alpha_{\text{me}}=\alpha_{\text{me}}^{\text{out}}=2\frac{e^{2}}{hc}\frac{N-1}{2N}+\frac{d_{s}}{2N(d_{s}+d_{v})}(\mathcal{A}_{+}^{2}+\mathcal{A}_{-}^{2}-\mathcal{A}_{+}^{2}\phi_{+}^{2}-\mathcal{A}_{-}^{2}\phi_{-}^{2}),$ (S26)
      which is $e^{2}/hc$ in the limit of $N\to\infty$.
    3. (c)
      $\delta>(\Delta_{\text{S}}/J_{\text{S}})^{2}$
      In this case, $(1-\delta/2)J_{\text{S}}-\sqrt{\Delta_{\text{S}}^{2}+(J_{\text{S}}\delta/2)^{2}}<0$ and $(1-\delta)J_{\text{S}}-\Delta_{\text{S}}<0$, the occupation of the states is exactly the same as the case of $J_{\text{S}}<\Delta_{\text{S}}$, and the magnetoelectric response goes to 0 in the thick limit of thin films.
For general values of $\delta$ and $\Delta_{\text{D}}$, the magnetoelectric coefficients calculated from numerical methods are shown in Fig. S1 (b)-(d), and the corresponding phase is shown in (a).
Appendix S2 Bulk calculations for three-dimensional MBX-type crystalline insulators
S2.1 A tight binding model
We employ a previously developed [15] square lattice regularization for the effective low-energy coupled Dirac cone model that was introduced (Eq. (2)) in the main text. The family of materials we consider are layered compounds (with stacking direction taken to be $\bm{z}$ and interlayer spacing $d$), and the tight-binding model we use was obtained by developing a 2D square lattice regularization of the $\bm{k}\cdot\bm{p}$ model of an isolated layer then coupling the states associated with the nearest surfaces of adjacent layers. In the non-magnetic case, the model is periodic under translations by $d\bm{z}$ (i.e., one layer per unit cell), while in the AFM case the model is periodic under translations by $2d\bm{z}$ (i.e., two layers per unit cell). The 3D square lattice regularized model (with Bravais lattice $\Gamma=\mathbb{Z}\times\mathbb{Z}\times 2d\mathbb{Z}$ and whose dual lattice is $\Gamma^{*}=2\pi\mathbb{Z}\times 2\pi\mathbb{Z}\times\frac{2\pi}{2d}\mathbb{Z}$) can be written in the general form
$\hat{H}^{(\text{3D})}_{\text{reg}}=\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\hat{c}^{\dagger}(\bm{k})\mathcal{H}^{(\text{3D})}_{\text{reg}}(\bm{k})\hat{c}(\bm{k}),$ (S27)
where $\mathcal{H}^{(\text{3D})}_{\text{reg}}(\bm{k})$ is a smooth, Hermitian-matrix-valued map in $\mathbb{R}^{3}$ that is $\Gamma^{*}$-periodic (i.e., $\forall\bm{G}\in\Gamma^{*}:\mathcal{H}^{(\text{3D})}_{\text{reg}}(\bm{k})=\mathcal{H}^{(\text{3D})}_{\text{reg}}(\bm{k}+\bm{G})$ is satisfied), BZ denotes a 3D Brillouin zone (BZ) of $\Gamma^{*}$, and
$\hat{c}(\bm{k})$ $\equiv\left(\hat{c}_{(0,\uparrow),\bm{k}},\hat{c}_{(0,\downarrow),\bm{k}},\hat{c}_{(1,\uparrow),\bm{k}},\hat{c}_{(1,\downarrow),\bm{k}},\hat{c}_{(2,\uparrow),\bm{k}},\hat{c}_{(2,\downarrow),\bm{k}},\hat{c}_{(3,\uparrow),\bm{k}},\hat{c}_{(3,\downarrow),\bm{k}}\right)^{\text{T}},$
$\hat{c}^{\dagger}(\bm{k})$ $\equiv\left(\hat{c}^{\dagger}_{(0,\uparrow),\bm{k}},\hat{c}^{\dagger}_{(0,\downarrow),\bm{k}},\hat{c}^{\dagger}_{(1,\uparrow),\bm{k}},\hat{c}^{\dagger}_{(1,\downarrow),\bm{k}},\hat{c}^{\dagger}_{(2,\uparrow),\bm{k}},\hat{c}^{\dagger}_{(2,\downarrow),\bm{k}},\hat{c}^{\dagger}_{(3,\uparrow),\bm{k}},\hat{c}^{\dagger}_{(3,\downarrow),\bm{k}}\right).$ (S28)
The operators $\hat{c}_{(\alpha,\sigma),\bm{k}}$, $\hat{c}^{\dagger}_{(\alpha,\sigma),\bm{k}}$ correspond to Block-type vectors |$\ket{\bar{\psi}_{(\alpha,\sigma),\bm{k}}}\equiv\hat{c}^{\dagger}_{(\alpha,\sigma),\bm{k}}\ket{\text{vac}}$ (i.e. they satisfy |$\ket{\bar{\psi}_{(\alpha,\sigma),\bm{k}+\bm{G}}}=\ket{\bar{\psi}_{(\alpha,\sigma),\bm{k}}}$ for any $\bm{G}\in\Gamma^{*}$) that are smooth over BZ, are mutually orthogonal, and map to the exponentially localized Wannier functions (WFs) |$\ket{W_{(\alpha,\sigma),\bm{R}}}$ [61] with respect to which the lattice regularized tight-binding model is written. We use even (odd) values of $\alpha$ to label WFs that are associated with the bottom (top) surface [62, 63] of each layer. If there are $L$ layers in a unit cell then the WFs |$\ket{W_{(\alpha,\sigma),\bm{R}}}$ associated with layer $l\in\{1,\ldots,L\}$ (enumerated from bottom layer to top layer) are labeled by $\alpha=2(l-1)$ and $12l-$ 1. We denote the corresponding cell-periodic parts of the Bloch-type vectors |$\ket{\bar{\psi}_{(\alpha,\sigma),\bm{k}}}$ by |$\ket{\bar{u}_{(\alpha,\sigma),\bm{k}}}$.
As previously described [15], the |$\ket{W_{(\alpha,\sigma),\bm{R}}}$ can be assumed to be atomic-like (this is almost always done without mention in tight-binding calculations), a characteristic property of which is $\partial_{a}\ket{\bar{u}_{(\alpha,\sigma),\bm{k}}}\equiv 0$ where $\partial_{a}\equiv\partial/\partial k^{a}$ and $a=1$, $2$, or $3$ ($\equiv x$, $y$, or $z$). A more physical (although still highly simplified) approximation is reached by assuming that each WF $W_{(i,\alpha),\bm{R}}(\bm{r})$ is extremely localized about a point $\bm{\delta}_{\alpha}$ in the unit cell at $\bm{R}$ with which it is associated, from which it follows [64]
$\frac{\partial}{\partial k^{a}}\bar{u}_{(\alpha,\sigma),\bm{k}}(\bm{r})=\sqrt{\frac{\Omega_{uc}}{(2\pi)^{3}}}\sum_{\bm{R}}\frac{\partial e^{-i\bm{k}\cdot(\bm{r}-\bm{R})}}{\partial k^{a}}W_{(\alpha,\sigma),\bm{R}}(\bm{r})\approx\sqrt{\frac{\Omega_{uc}}{(2\pi)^{3}}}\sum_{\bm{R}}\frac{\partial e^{-i\bm{k}\cdot\bm{\delta}_{\alpha}}}{\partial k^{a}}W_{(\alpha,\sigma),\bm{R}}(\bm{r})=-i\delta_{\alpha}^{a}\bar{u}_{(\alpha,\sigma),\bm{k}}(\bm{r}).$ (S29)
In the non-magnetic case, there is one layer per unit cell and if the layer thickness is neglected then one can choose $\bm{\delta}_{\alpha}=\bm{0}$, which is equivalent to the assumption of atomic-like WFs. However, in the AFM case there are two layers per unit cell, thus there is always more than one position within it about which WFs are localized. In particular, even if the WFs within each layer are taken to be localized about the same point, $\bm{\delta}_{0}=\bm{\delta}_{1}$ and $\bm{\delta}_{2}=\bm{\delta}_{3}$, we have $\bm{\delta}_{2}-\bm{\delta}_{0}=d\bm{z}$. In this approximation, the most natural choice is to take $\bm{\delta}_{2}=\frac{d}{2}\bm{z}$ and $\bm{\delta}_{0}=-\frac{d}{2}\bm{z}$.
Under our lattice regularization scheme, the general AFM Hamiltonian matrix is
$\mathcal{H}^{(\text{3D})}_{\text{reg}}(\bm{k})=$
$\left(\begin{array}[]{cccccccc}\mu&iA(\text{s}k_{x}-i\text{s}k_{y})&\Delta_{\text{S}}(k_{x},k_{y})&0&0&0&e^{-2idk_{z}}\Delta_{\text{D}}&0\\ -iA(\text{s}k_{x}+i\text{s}k_{y})&-\mu&0&\Delta_{\text{S}}(k_{x},k_{y})&0&0&0&e^{-2idk_{z}}\Delta_{\text{D}}\\ \Delta_{\text{S}}(k_{x},k_{y})&0&\mu&-iA(\text{s}k_{x}-i\text{s}k_{y})&\Delta_{\text{D}}&0&0&0\\ 0&\Delta_{\text{S}}(k_{x},k_{y})&iA(\text{s}k_{x}+i\text{s}k_{y})&-\mu&0&\Delta_{\text{D}}&0&0\\ 0&0&\Delta_{\text{D}}&0&-\mu&iA(\text{s}k_{x}-i\text{s}k_{y})&\Delta_{\text{S}}(k_{x},k_{y})&0\\ 0&0&0&\Delta_{\text{D}}&-iA(\text{s}k_{x}+i\text{s}k_{y})&\mu&0&\Delta_{\text{S}}(k_{x},k_{y})\\ e^{2idk_{z}}\Delta_{\text{D}}&0&0&0&\Delta_{\text{S}}(k_{x},k_{y})&0&-\mu&-iA(\text{s}k_{x}-i\text{s}k_{y})\\ 0&e^{2idk_{z}}\Delta_{\text{D}}&0&0&0&\Delta_{\text{S}}(k_{x},k_{y})&iA(\text{s}k_{x}+i\text{s}k_{y})&\mu\end{array}\right).$ (S38)
Here $\Delta_{\text{S}}(k_{x},k_{y})\equiv\Delta_{\text{S}}-2B(2-\text{c}k_{x}-\text{c}k_{y})$, where $\text{c}k_{x}\equiv\cos(k_{x})$, $\text{s}k_{x}\equiv\sin(k_{x})$, etc. This Hamiltonian is valid only when $\mu\neq 0$, since otherwise the material is non-magnetic and the crystalline unit cell is misidentified [15]. The eigenvalues of (S38) are
$E_{1,2}(\bm{k})$ $=-E_{7,8}(\bm{k})=-\sqrt{\frac{1}{2}A^{2}(2-\text{c}2k_{x}-\text{c}2k_{y})+\Delta_{\text{S}}(k_{x},k_{y})^{2}+\Delta_{\text{D}}^{2}+\mu^{2}+2\sqrt{\Delta_{\text{S}}(k_{x},k_{y})^{2}(\Delta_{\text{D}}^{2}\cos^{2}(dk_{z})+\mu^{2})}},$
$E_{3,4}(\bm{k})$ $=-E_{5,6}(\bm{k})=-\sqrt{\frac{1}{2}A^{2}(2-\text{c}2k_{x}-\text{c}2k_{y})+\Delta_{\text{S}}(k_{x},k_{y})^{2}+\Delta_{\text{D}}^{2}+\mu^{2}-2\sqrt{\Delta_{\text{S}}(k_{x},k_{y})^{2}(\Delta_{\text{D}}^{2}\cos^{2}(dk_{z})+\mu^{2})}}.$ (S39)
We will consider the ground state of the material to be that for which the energy bands are half-filled. We restrict our focus to band insulators, which implies $\Delta_{\text{S}}^{2}+\Delta_{\text{D}}^{2}+\mu^{2}-2\sqrt{\Delta_{\text{S}}^{2}(\Delta_{\text{D}}^{2}+\mu^{2})}>0$ (i.e., $|\Delta_{\text{S}}|-\sqrt{\Delta_{\text{D}}^{2}+\mu^{2}}\neq 0$). Moreover, the family of MBX materials of interest are such that the band gap is smallest at $\bm{k}=(0,0,0)$, which is always the case when $B/\Delta_{\text{S}}<0$. The double degeneracy of the energy bands at each $\bm{k}\in\text{BZ}$ follows from the combination of an inversion and a (fermionic) time-reversal symmetry (see Appendix C of Ref. [15]). A consequence is that the corresponding eigenvectors are highly non-unique. We wish to identify a set of orthogonal energy eigenvectors that are smooth and orthogonal over BZ such that they constitute a smooth global (and periodic) gauge choice of the bundle of occupied Bloch states over the Brillouin zone torus and of the total Bloch bundle more generally. One such choice, written in the basis (|$\big{(}\ket{\bar{\psi}_{(\alpha,\uparrow),\bm{k}}},\,\ket{\bar{\psi}_{(\alpha,\downarrow),\bm{k}}}\big{)}_{\alpha\in\{0,1,2,3\}}$, is
$\Psi_{1,\bm{k}}$ $=\Bigg{(}-\frac{(1+e^{2idk_{z}})\Delta_{\text{S}}(k_{x},k_{y})+\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}}{(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{1}(\bm{k})},\,0,\,1,\,\frac{(1+e^{2idk_{z}})A(-i\text{s}k_{x}+\text{s}k_{y})}{(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{1}(\bm{k})},$
$\qquad-\frac{(1+e^{2idk_{z}})\Delta_{\text{D}}^{2}+e^{2idk_{z}}\Big{(}2\mu(\mu-E_{1}(\bm{k}))+\sqrt{\Delta_{\text{S}}(k_{x},k_{y})^{2}(2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z})))}\Big{)}}{\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{1}(\bm{k}))},$
$\qquad\frac{2e^{2idk_{z}}A\mu(-i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{1}(\bm{k})\big{)}},\,\frac{2e^{2idk_{z}}\big{(}\Delta_{\text{S}}(k_{x},k_{y})\mu+\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(\mu-E_{1}(\bm{k}))\big{)}}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{1}(\bm{k})\big{)}},$
$\qquad\frac{2e^{2idk_{z}}A\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(-i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{1}(\bm{k})\big{)}}\Bigg{)},$
$\Psi_{7,\bm{k}}$ $=\Bigg{(}-\frac{(1+e^{2idk_{z}})\Delta_{\text{S}}(k_{x},k_{y})+\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}}{(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{7}(\bm{k})},\,0,\,1,\,\frac{(1+e^{2idk_{z}})A(-i\text{s}k_{x}+\text{s}k_{y})}{(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{7}(\bm{k})},$
$\qquad-\frac{(1+e^{2idk_{z}})\Delta_{\text{D}}^{2}+e^{2idk_{z}}\Big{(}2\mu(\mu-E_{7}(\bm{k}))+\sqrt{\Delta_{\text{S}}(k_{x},k_{y})^{2}(2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z})))}\Big{)}}{\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{7}(\bm{k}))},$
$\qquad\frac{2e^{2idk_{z}}A\mu(-i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{7}(\bm{k})\big{)}},\,\frac{2e^{2idk_{z}}\big{(}\Delta_{\text{S}}(k_{x},k_{y})\mu+\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(\mu-E_{7}(\bm{k}))\big{)}}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{7}(\bm{k})\big{)}},$
$\qquad\frac{2e^{2idk_{z}}A\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(-i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{7}(\bm{k})\big{)}}\Bigg{)},$ (S40a)
$\Psi_{3,\bm{k}}$ $=\Bigg{(}\frac{-(1+e^{2idk_{z}})\Delta_{\text{S}}(k_{x},k_{y})+\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}}{(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{3}(\bm{k})},\,0,\,1,\,\frac{(1+e^{2idk_{z}})A(-i\text{s}k_{x}+\text{s}k_{y})}{(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{3}(\bm{k})},$
$\qquad-\frac{(1+e^{2idk_{z}})\Delta_{\text{D}}^{2}+e^{2idk_{z}}\Big{(}2\mu(\mu-E_{3}(\bm{k}))-\sqrt{\Delta_{\text{S}}(k_{x},k_{y})^{2}(2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z})))}\Big{)}}{\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{3}(\bm{k}))},$
$\qquad\frac{2e^{2idk_{z}}A\mu(-i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{3}(\bm{k})\big{)}},\,\frac{2e^{2idk_{z}}\big{(}\Delta_{\text{S}}(k_{x},k_{y})\mu-\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(\mu-E_{3}(\bm{k}))\big{)}}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{3}(\bm{k})\big{)}},$
$\qquad-\frac{2e^{2idk_{z}}A\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(-i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{3}(\bm{k})\big{)}}\Bigg{)},$
$\Psi_{5,\bm{k}}$ $=\Bigg{(}\frac{-(1+e^{2idk_{z}})\Delta_{\text{S}}(k_{x},k_{y})+\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}}{(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{5}(\bm{k})},\,0,\,1,\,\frac{(1+e^{2idk_{z}})A(-i\text{s}k_{x}+\text{s}k_{y})}{(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{5}(\bm{k})},$
$\qquad-\frac{(1+e^{2idk_{z}})\Delta_{\text{D}}^{2}+e^{2idk_{z}}\Big{(}2\mu(\mu-E_{5}(\bm{k}))-\sqrt{\Delta_{\text{S}}(k_{x},k_{y})^{2}(2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z})))}\Big{)}}{\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{5}(\bm{k}))},$
$\qquad\frac{2e^{2idk_{z}}A\mu(-i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{5}(\bm{k})\big{)}},\,\frac{2e^{2idk_{z}}\big{(}\Delta_{\text{S}}(k_{x},k_{y})\mu-\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(\mu-E_{5}(\bm{k}))\big{)}}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{5}(\bm{k})\big{)}},$
$\qquad-\frac{2e^{2idk_{z}}A\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(-i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{5}(\bm{k})\big{)}}\Bigg{)},$ (S40b)
$\Psi_{2,\bm{k}}$ $=\Bigg{(}\frac{A\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}(i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{2}(\bm{k})\big{)}},\,-\frac{2\Delta_{\text{s}}(k_{x},k_{y})\mu+2\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(\mu-E_{2}(\bm{k}))}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{2}(\bm{k})\big{)}},$
$\frac{2A\mu(i\text{s}(k_{x})+\text{s}(k_{y}))}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{2}(\bm{k})\big{)}},$
$\frac{((-1+e^{2idk_{z}})\Delta_{\text{D}}^{2}\mu-2A^{2}\mu(\text{s}k_{x}^{2}+\text{s}k_{y}^{2})-2\mu\Delta_{\text{S}}(k_{x},k_{y})^{2}+e^{idk_{z}}\Delta_{\text{D}}^{2}\text{c}(dk_{z})2E_{2}(\bm{k})+\sqrt{(\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2})\Delta_{\text{S}}(k_{x},k_{y})^{2}}(-2\mu+2E_{2}(\bm{k})))}{(\mu+E_{2}(\bm{k}))\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{2}(\bm{k}))},$
$-\frac{(1+e^{2idk_{z}})A(-i\text{s}k_{x}-\text{s}k_{y})}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{2}(\bm{k})},\,1,\,0,\,\frac{(1+e^{2idk_{z}})\Delta_{\text{s}}(k_{x},k_{y})+e^{2idk_{z}}\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{2}(\bm{k})}\Bigg{)},$
$\Psi_{8,\bm{k}}$ $=\Bigg{(}\frac{A\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}(i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{8}(\bm{k})\big{)}},\,-\frac{2\Delta_{\text{s}}(k_{x},k_{y})\mu+2\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(\mu-E_{8}(\bm{k}))}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{8}(\bm{k})\big{)}},$
$\frac{2A\mu(i\text{s}(k_{x})+\text{s}(k_{y}))}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{8}(\bm{k})\big{)}},$
$\frac{((-1+e^{2idk_{z}})\Delta_{\text{D}}^{2}\mu-2A^{2}\mu(\text{s}k_{x}^{8}+\text{s}k_{y}^{8})-2\mu\Delta_{\text{S}}(k_{x},k_{y})^{2}+e^{idk_{z}}\Delta_{\text{D}}^{2}\text{c}(dk_{z})2E_{8}(\bm{k})+\sqrt{(\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2})\Delta_{\text{S}}(k_{x},k_{y})^{2}}(-2\mu+2E_{8}(\bm{k})))}{(\mu+E_{8}(\bm{k}))\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{8}(\bm{k}))},$
$-\frac{(1+e^{2idk_{z}})A(-i\text{s}k_{x}-\text{s}k_{y})}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{8}(\bm{k})},\,1,\,0,\,\frac{(1+e^{2idk_{z}})\Delta_{\text{s}}(k_{x},k_{y})+e^{2idk_{z}}\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{8}(\bm{k})}\Bigg{)},$ (S41a)
$\Psi_{4,\bm{k}}$ $=\Bigg{(}-\frac{A\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}(i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{4}(\bm{k})\big{)}},\,-\frac{2\Delta_{\text{s}}(k_{x},k_{y})\mu-2\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(\mu-E_{4}(\bm{k}))}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{4}(\bm{k})\big{)}},$
$\frac{2A\mu(i\text{s}(k_{x})+\text{s}(k_{y}))}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{4}(\bm{k})\big{)}},$
$\frac{((-1+e^{2idk_{z}})\Delta_{\text{D}}^{2}\mu-2A^{2}\mu(\text{s}k_{x}^{2}+\text{s}k_{y}^{2})-2\mu\Delta_{\text{S}}(k_{x},k_{y})^{2}+e^{idk_{z}}\Delta_{\text{D}}^{2}\text{c}(dk_{z})2E_{4}(\bm{k})-\sqrt{(\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2})\Delta_{\text{S}}(k_{x},k_{y})^{2}}(-2\mu+2E_{4}(\bm{k})))}{(\mu+E_{4}(\bm{k}))\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{4}(\bm{k}))},$
$-\frac{(1+e^{2idk_{z}})A(-i\text{s}k_{x}-\text{s}k_{y})}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{4}(\bm{k})},\,1,\,0,\,\frac{(1+e^{2idk_{z}})\Delta_{\text{s}}(k_{x},k_{y})-e^{2idk_{z}}\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{4}(\bm{k})}\Bigg{)},$
$\Psi_{6,\bm{k}}$ $=\Bigg{(}-\frac{A\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}(i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{6}(\bm{k})\big{)}},\,-\frac{2\Delta_{\text{s}}(k_{x},k_{y})\mu-2\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(\mu-E_{6}(\bm{k}))}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{6}(\bm{k})\big{)}},$
$\frac{2A\mu(i\text{s}(k_{x})+\text{s}(k_{y}))}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{6}(\bm{k})\big{)}},$
$\frac{((-1+e^{2idk_{z}})\Delta_{\text{D}}^{2}\mu-2A^{2}\mu(\text{s}k_{x}^{2}+\text{s}k_{y}^{2})-2\mu\Delta_{\text{S}}(k_{x},k_{y})^{2}+e^{idk_{z}}\Delta_{\text{D}}^{2}\text{c}(dk_{z})2E_{6}(\bm{k})-\sqrt{(\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2})\Delta_{\text{S}}(k_{x},k_{y})^{2}}(-2\mu+2E_{6}(\bm{k})))}{(\mu+E_{6}(\bm{k}))\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{6}(\bm{k}))},$
$-\frac{(1+e^{2idk_{z}})A(-i\text{s}k_{x}-\text{s}k_{y})}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{6}(\bm{k})},\,1,\,0,\,\frac{(1+e^{2idk_{z}})\Delta_{\text{s}}(k_{x},k_{y})-e^{2idk_{z}}\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{6}(\bm{k})}\Bigg{)}.$ (S41b)
At least one component of each of the $\Psi_{n,\bm{k}}$ in (S40,S41) is nonzero at every $\bm{k}\in\text{BZ}$, thus $\forall\bm{k}\in\text{BZ}:\innerproduct{\Psi_{n,\bm{k}}}{\Psi_{n,\bm{k}}}\neq 0$. Indeed for every $\bm{k}\in\text{BZ}$ each of the vectors |$\ket{\Psi_{1,\bm{k}}},\ldots,\ket{\Psi_{8,\bm{k}}}$ defined by Eq. (S40,S41) are energy eigenvectors of Bloch’s form (they satisfy |$\ket{\Psi_{n,\bm{k}+\bm{G}}}=\ket{\Psi_{n,\bm{k}}}$) and are smooth over BZ. These vectors can be normalized to obtain
|$\ket{\psi_{n,\bm{k}}}\equiv\frac{\ket{\Psi_{n,\bm{k}}}}{\innerproduct{\Psi_{n,\bm{k}}}{\Psi_{n,\bm{k}}}}.$ (S42)
We denote the corresponding cell-periodic functions by |$\ket{u_{n,\bm{k}}}$. Thus we identify a smooth global Hamiltonian gauge $\mathfrak{u}$ pointwise by $\mathfrak{u}_{\bm{k}}=(\ket{u_{n,\bm{k}}})_{n\in\{1,\ldots,8\}}$.
We identify the components of the non-Abelian Berry connection induced by the gauge $\mathfrak{u}$ by
$\xi^{a}_{nm}(\bm{k})\equiv\big{(}\xi^{\mathfrak{u}}(\bm{k})\big{)}^{a}_{nm}=i\innerproduct{u_{n,\bm{k}}}{\partial_{a}u_{m,\bm{k}}}.$ (S43)
In principle, the components (S43) can be obtained by explicitly taking $\bm{k}$-derivatives of Eqs. (S40,S41) and employing some approximation for the $\bm{k}$-derivatives of the |$\ket{\bar{u}_{(\alpha,\sigma),\bm{k}}}$ (for example, using the assumption of atomic-like WFs or of highly-localized WFs as described above). However, we will perform our calculations numerically. Some useful (numerically obtained) relations that arise in this gauge choice are
$\xi^{a}_{11}=-\xi^{a}_{22},\,\xi^{a}_{33}=-\xi^{a}_{44},\,\xi^{a}_{55}=-\xi^{a}_{66},\,\xi^{a}_{77}=-\xi^{a}_{88},$
$\xi^{a}_{13}=-\xi^{a}_{42},\,\xi^{a}_{14}=\xi^{a}_{32},\,\xi^{a}_{57}=-\xi^{a}_{86},\,\xi^{a}_{58}=\xi^{a}_{76},$
$\xi^{a}_{15}=-\xi^{a}_{62},\,\xi^{a}_{16}=\xi^{a}_{52},\,\xi^{a}_{17}=-\xi^{a}_{82},\,\xi^{a}_{18}=\xi^{a}_{72},$
$\xi^{a}_{35}=-\xi^{a}_{64},\,\xi^{a}_{36}=\xi^{a}_{54},\,\xi^{a}_{37}=-\xi^{a}_{84},\,\xi^{a}_{38}=\xi^{a}_{74}.$ (S44)
The relations (S44) arise in both the approximation of atomic-like WFs ($\partial_{a}\ket{\bar{u}_{(\alpha,\sigma),\bm{k}}}\equiv 0$) and the approximation of highly-localized WFs (S29) for $\bm{\delta}_{2}=\bm{\delta}_{3}=\frac{d}{2}\bm{z}$ and $\bm{\delta}_{0}=\bm{\delta}_{1}=-\frac{d}{2}\bm{z}$. These relations are analogous to those presented in Eq. (42) of Ref. [15] for the non-magnetic case. Indeed each of the relations in Eq. (S44) have a natural analog in the non-magnetic case. However, in the non-magnetic case there are relations that have no analog here; in the Hamiltonian gauge employed in the non-magnetic case [15], $\xi^{a}_{vv^{\prime}}\propto\xi^{a}_{cc^{\prime}}$ for various combinations of $v$, $v^{\prime}$, $c$, and $c^{\prime}$ (in particular, $\xi^{a}_{11}=\xi^{a}_{33}$, $\xi^{a}_{22}=\xi^{a}_{44}$ and $\xi^{a}_{12}=-\xi^{a}_{34}$, where bands 1 and 2 are filled and bands 3 and 4 are empty). In the AFM case we do not have a relation between any $\xi^{a}_{vv}$ and $\xi^{a}_{cc}$, nor do we have a relation between $\xi^{a}_{12}$ and $\xi^{a}_{78}$ or between $\xi^{a}_{34}$ and $\xi^{a}_{56}$. As we later demonstrate, the lack of such relations will result in a qualitative distinction between the partitioning of the magnetoelectric response into atomic-like and itinerant contributions in the non-magnetic and AFM cases.
In the AFM case there also emerges a property of certain $\xi^{a}_{nm}$ that is not present in the non-magnetic case. It turns out that certain band components of the AFM $\xi^{a}$ are numerically non-negligible only when $a=z$. In particular,
$\xi^{a}_{13},\,\xi^{a}_{14},\,\xi^{a}_{15},\,\xi^{a}_{16},\,\xi^{a}_{23},\,\xi^{a}_{24},\,\xi^{a}_{25},\,\xi^{a}_{26},\,\xi^{a}_{73},\,\xi^{a}_{74},\,\xi^{a}_{75},\,\xi^{a}_{76},\,\xi^{a}_{83},\,\xi^{a}_{84},\,\xi^{a}_{85},\,\xi^{a}_{86}$ $\neq 0\text{ only if }a=3.$ (S45)
S2.2 Computation of $\alpha_{\text{CS}}$
We can now directly compute the topological magnetoelectric coefficient in the gauge $\mathfrak{u}$, which is given by
$\alpha^{\mathfrak{u}}_{\text{CS}}=-\frac{e^{2}}{2\hbar c}\epsilon^{abd}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\left(\sum_{vv^{\prime}}\xi^{a}_{vv^{\prime}}(\bm{k})\partial_{b}\xi^{d}_{v^{\prime}v}(\bm{k})-\frac{2i}{3}\sum_{vv^{\prime}v_{1}}\xi^{a}_{vv^{\prime}}(\bm{k})\xi^{b}_{v^{\prime}v_{1}}(\bm{k})\xi^{d}_{v_{1}v}(\bm{k})\right),$ (S46)
where $v,v^{\prime},v_{1}\in\{1,2,3,4\}$ in this case. Again, the gauge $\mathfrak{u}$ is defined by Eq. (S40,S41). We evaluate this integral numerically and interpolate to convergence. Some results are listed in Table 1. To correspond with MBX materials we focus on the parameter regime $B/\Delta_{\text{S}}<0$, and arbitrarily choose $\Delta_{\text{S}}=190$ meV and $B=-25$ meV. The values of $\alpha^{\mathfrak{u}}_{\text{CS}}$ listed in Table 1 are independent of $A\in\mathbb{R}_{+}$, which is a parameter that is related to the single layer lattice constant that was introduced during regularization of the effective low-energy coupled Dirac cone model. Notably, when the values of the crystal parameters are varied such that the pairs of degenerate bands remain isolated from one another, the value to which $\alpha^{\mathfrak{u}}_{\text{CS}}$ evaluates is unchanged. The findings listed in Table 1 are consistent with Fig. 2 (a) of the main text. Due to the large number of parameters in our model (Eq. (S38)) we do not extract the entire topological phase diagram. Via explicit calculation we have found that the value to which $\alpha^{\mathfrak{u}}_{\text{CS}}$ evaluates is unchanged regardless of whether we adopt either of the above described approximations of atomic-like WFs ($\partial_{a}\ket{\bar{u}_{(\alpha,\sigma),\bm{k}}}\equiv 0$) or of highly-localized WFs (S29) with $\bm{\delta}_{2}=\bm{\delta}_{3}=\frac{d}{2}\bm{z}$ and $\bm{\delta}_{0}=\bm{\delta}_{1}=-\frac{d}{2}\bm{z}$.
$J_{\text{S}}/\Delta_{\text{S}}$ $\Delta_{\text{D}}/\Delta_{\text{S}}$ $J_{\text{D}}/J_{\text{S}}$ $\alpha^{\mathfrak{u}}_{\text{CS}}$ $(\frac{e^{2}}{hc})$
0.8 ±$\pm 0.25$
0.25
0.5
0
0
0.8 ±$\pm 1.25$
0.25
0.5
1/2
1/2
   $J_{\text{S}}/\Delta_{\text{S}}$ $\Delta_{\text{D}}/\Delta_{\text{S}}$ $J_{\text{D}}/J_{\text{S}}$ $\alpha^{\mathfrak{u}}_{\text{CS}}$ $(\frac{e^{2}}{hc})$ 1.5 ±$\pm 0.25$ 0.25 0.5 1/2 0 1.5 ±$\pm 1.25$ 0.25 0.5 -1/2 3/2
Table 1: Values of $\alpha^{\mathfrak{u}}_{\text{CS}}$ obtained from numerical integration of Eq. (S46) for $\Delta_{\text{S}}=190$ meV and $B=-25$ meV.
S2.3 Partitioning of $\alpha_{\text{CS}}$ into atomic-like and itinerant contributions
S2.3.1 Considerations for a general smooth global Hamiltonian gauge
We begin with a discussion of the linear response of orbital magnetization in a band insulator that exhibits TRS. We then restrict focus to our model, Eq. (S38). In general, the electronic orbital magnetization $\bm{M}$ can be partitioned into a sum of atomic-like $\overline{\bm{M}}$ and itinerant $\widetilde{\bm{M}}$ contributions,
$\bm{M}=\overline{\bm{M}}+\widetilde{\bm{M}}.$ (S47)
This decomposition applies both in the ground state, $\bm{M}^{(0)}=\overline{\bm{M}}^{(0)}+\widetilde{\bm{M}}^{(0)}$ [47, 48, 65], and at linear response to an electric field, $\bm{M}^{(E)}=\overline{\bm{M}}^{(E)}+\widetilde{\bm{M}}^{(E)}$ [66, 67]. We previously [15] expounded on this decomposition. We now employ expressions that were previously derived [67] for the linear response of these contributions, and also use the results presented there that identify the terms in $\overline{\bm{M}}^{(E)}$ and $\widetilde{\bm{M}}^{(E)}$ that give rise to $\alpha_{\text{CS}}$ and $\alpha^{il}_{\text{G}}$.
We employ Eq. (70) and (71,72) of Ref. [67] for $\overline{\bm{M}}^{(E)}$ and $\widetilde{\bm{M}}^{(E)}$, respectively, which applies to any crystalline insulator whose electronic ground state is characterized by a vanishing (triple of) Chern number(s). We will initially work in a general smooth (and periodic) Hamiltonian gauge, such that $U_{n\alpha}(\bm{k})\equiv\delta_{n\alpha}$ in those expressions; the WFs with respect to which the electronic contribution to the electric polarization and orbital magnetization are identified are |$\ket{W_{n,\bm{R}}}\propto\int_{\text{BZ}}d^{3}ke^{-i\bm{k}\cdot\bm{R}}\ket{\phi_{n,\bm{k}}}$. That is, we initially consider our gauge choice to be defined by unspecified Bloch energy eigenvectors |$\ket{\phi_{n,\bm{k}}}$ that are smooth over BZ and only later will we take |$\ket{\phi_{n,\bm{k}}}=\ket{\psi_{n,\bm{k}}}$ defined by Eq. (S40,S41). We denote the corresponding cell-periodic functions by |$\ket{v_{n,\bm{k}}}$ and identify a smooth global Hamiltonian gauge $\mathfrak{v}$ pointwise by $\mathfrak{v}_{\bm{k}}=(\ket{v_{n,\bm{k}}})_{n\in\{1,\ldots,N\}}$ where $N$ is the number of energy bands in the model. We choose to work in a Hamiltonian gauge because we seek the existence of a smooth global gauge in which Eq. (70) of Ref. [67] vanishes, and in any Hamiltonian gauge the terms involving $\mathcal{W}^{a}$ vanish. Otherwise, those terms appear to be generically nonzero. Perhaps in special cases smooth global symmetric gauge choices exist (although in a $\mathbb{Z}_{2}$-odd phase a time-reversal symmetric smooth global frame is topologically forbidden [68, 44]) for which these additional terms vanish. Nevertheless, we focus on smooth global Hamiltonian gauge choices. In such a gauge, Eq. (70) of Ref. [67] reduces to
$\overline{M}_{\mathfrak{v}}^{i(E)}$ $=\frac{e^{2}}{4\hbar c}\epsilon^{iab}E^{l}\sum_{nm}f_{nm}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\Bigg{(}\frac{\partial_{b}(E_{n,\bm{k}}+E_{m,\bm{k}})}{E_{m,\bm{k}}-E_{n,\bm{k}}}(\xi^{\mathfrak{v}})^{a}_{nm}(\xi^{\mathfrak{v}})^{l}_{mn}+2\sum_{s}\frac{E_{s,\bm{k}}-E_{m,\bm{k}}}{E_{m,\bm{k}}-E_{n,\bm{k}}}\text{Re}\big{[}i(\xi^{\mathfrak{v}})^{a}_{ns}(\xi^{\mathfrak{v}})^{b}_{sm}(\xi^{\mathfrak{v}})^{l}_{mn}\big{]}\Bigg{)}$
$=\frac{e^{2}}{4\hbar c}\epsilon^{iab}E^{l}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\Bigg{(}\sum_{nm}f_{nm}\frac{\partial_{b}(E_{n,\bm{k}}+E_{m,\bm{k}})}{E_{m,\bm{k}}-E_{n,\bm{k}}}(\xi^{\mathfrak{v}})^{a}_{nm}(\xi^{\mathfrak{v}})^{l}_{mn}+4\sum_{vv^{\prime}c}\frac{E_{v^{\prime},\bm{k}}-E_{v,\bm{k}}}{E_{c,\bm{k}}-E_{v,\bm{k}}}\text{Re}\big{[}i(\xi^{\mathfrak{v}})^{a}_{vv^{\prime}}(\xi^{\mathfrak{v}})^{b}_{v^{\prime}c}(\xi^{\mathfrak{v}})^{l}_{cv}\big{]}$
$\qquad\qquad\qquad\qquad\qquad\qquad-4\sum_{vcc^{\prime}}\frac{E_{c^{\prime},\bm{k}}-E_{c,\bm{k}}}{E_{v,\bm{k}}-E_{c,\bm{k}}}\text{Re}\big{[}i(\xi^{\mathfrak{v}})^{a}_{cc^{\prime}}(\xi^{\mathfrak{v}})^{b}_{c^{\prime}v}(\xi^{\mathfrak{v}})^{l}_{vc}\big{]}\Bigg{)}$
$+\frac{e^{2}}{2\hbar c}\epsilon^{iab}E^{l}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\text{Re}\Bigg{[}-2i\sum_{vv^{\prime}c}(\xi^{\mathfrak{v}})^{a}_{cv^{\prime}}(\xi^{\mathfrak{v}})^{b}_{v^{\prime}v}(\xi^{\mathfrak{v}})^{l}_{vc}+\sum_{vc}(\partial_{a}(\xi^{\mathfrak{v}})^{b}_{cv})(\xi^{\mathfrak{v}})^{l}_{vc}\Bigg{]},$ (S48)
where $\big{(}\xi^{\mathfrak{v}}\big{)}^{a}_{nm}(\bm{k})=i\innerproduct{v_{n,\bm{k}}}{\partial_{a}v_{m,\bm{k}}}$ are the components of the non-Abelian Berry connection induced by the gauge $\mathfrak{v}$. In what follows we will often not explicitly indicate the $\bm{k}$-dependence of quantities. Now recall that the explicit form of $\alpha^{il}_{\text{G}}$ is given by
$\alpha^{il}_{\text{G}}$ $=\frac{e^{2}}{4\hbar c}\epsilon^{iab}E^{l}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\Bigg{(}2\sum_{nm}f_{nm}\frac{\partial_{b}(E_{n,\bm{k}}+E_{m,\bm{k}})}{E_{m,\bm{k}}-E_{n,\bm{k}}}\xi^{a}_{nm}\xi^{l}_{mn}+4\sum_{vv^{\prime}c}\frac{E_{v^{\prime},\bm{k}}-E_{v,\bm{k}}}{E_{c,\bm{k}}-E_{v,\bm{k}}}\text{Re}\big{[}i\xi^{a}_{vv^{\prime}}\xi^{b}_{v^{\prime}c}\xi^{l}_{cv}\big{]}$
$\qquad\qquad\qquad\qquad\qquad\qquad-4\sum_{vcc^{\prime}}\frac{E_{c^{\prime},\bm{k}}-E_{c,\bm{k}}}{E_{v,\bm{k}}-E_{c,\bm{k}}}\text{Re}\big{[}i\xi^{a}_{cc^{\prime}}\xi^{b}_{c^{\prime}v}\xi^{l}_{vc}\big{]}\Bigg{)},$ (S49)
which is gauge invariant, hence the lack of a gauge label. In general, (generalized) TRS implies $\alpha^{il}_{\text{G}}=0$. In addition, note that in the second and third terms in the brackets of Eq. (S49), Berry connection matrix elements appear ($\xi^{a}_{vv^{\prime}}$ and $\xi^{a}_{cc^{\prime}}$, respectively) that can be independently varied by changing the crystal parameters of the Bloch Hamiltonian (while maintaining TRS). Thus, each of the terms in the brackets of Eq. (S49) must separately vanish if there is TRS. This implies that if the insulator exhibits TRS then all but the last line in the final equality in Eq. (S48) vanishes,
$\overline{M}_{\mathfrak{v}}^{i(E)}$ ==$\overset{\mathrm{TRS}}{=\joinrel=}\frac{e^{2}}{2\hbar c}\epsilon^{iab}E^{l}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\text{Re}\Bigg{[}-2i\sum_{vv^{\prime}c}(\xi^{\mathfrak{v}})^{a}_{cv^{\prime}}(\xi^{\mathfrak{v}})^{b}_{v^{\prime}v}(\xi^{\mathfrak{v}})^{l}_{vc}+\sum_{vc}(\partial_{a}(\xi^{\mathfrak{v}})^{b}_{cv})(\xi^{\mathfrak{v}})^{l}_{vc}\Bigg{]}$
$=\frac{e^{2}}{2\hbar c}\epsilon^{iab}E^{l}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\text{Re}\Bigg{[}-i\sum_{vv^{\prime}c}(\xi^{\mathfrak{v}})^{a}_{cv^{\prime}}(\xi^{\mathfrak{v}})^{b}_{v^{\prime}v}(\xi^{\mathfrak{v}})^{l}_{vc}+i\sum_{vcc^{\prime}}(\xi^{\mathfrak{v}})^{a}_{cc^{\prime}}(\xi^{\mathfrak{v}})^{b}_{c^{\prime}v}(\xi^{\mathfrak{v}})^{l}_{vc}\Bigg{]}$
$\equiv\bar{\alpha}_{\mathfrak{v}}^{li}E^{l}.$ (S50)
We can now implement any smooth global Hamiltonian gauge $\mathfrak{v}$ to evaluate Eq. (S50).
Consider any two gauge choices $\mathfrak{u}$ and $\mathfrak{v}$ of this type, which are identified pointwise as $\mathfrak{u}_{\bm{k}}=(\ket{u_{n,\bm{k}}})_{n\in\{1,\ldots,N\}}$ and $\mathfrak{v}_{\bm{k}}=(\ket{v_{n,\bm{k}}})_{n\in\{1,\ldots,N\}}$, respectively, where |$\ket{u_{n,\bm{k}}}$ and |$\ket{v_{n,\bm{k}}}$ are the cell-periodic part of Bloch energy eigenvectors |$\ket{\psi_{n,\bm{k}}}$ and |$\ket{\phi_{n,\bm{k}}}$. Any sets {|$\{\ket{\phi_{n,\bm{k}}}\}_{n,\bm{k}}$ and {|$\{\ket{\psi_{n,\bm{k}}}\}_{n,\bm{k}}$ of orthogonal energy eigenvectors are related at each $\bm{k}\in\text{BZ}$ by a $\Gamma^{*}$-periodic unitary transformation $T(\bm{k})$ and so too are their cell-periodic parts,
|$\ket{v_{n,\bm{k}}}=\sum_{m=1}^{N}\ket{u_{m,\bm{k}}}T_{mn}(\bm{k}),$ (S51)
where $T_{mn}(\bm{k})\equiv\big{(}T(\bm{k})\big{)}_{mn}$. In the case of a band insulator with $M\leq N$ occupied bands, $T(\bm{k})$ is represented by a block-diagonal unitary matrix with blocks of size $M\times M$ and $(N-M)\times(N-M)$. Then
$(\xi^{\mathfrak{v}})^{a}_{nm}(\bm{k})\equiv i\innerproduct{v_{n,\bm{k}}}{\partial_{a}v_{m,\bm{k}}}=\sum_{r,s=1}^{N}T^{\dagger}_{nr}(\bm{k})\big{(}(\xi^{\mathfrak{u}})^{a}_{rs}(\bm{k})+\mathcal{T}^{a}_{rs}(\bm{k})\big{)}T_{sm}(\bm{k}),$ (S52)
where $\mathcal{T}^{a}_{rs}(\bm{k})\equiv i\sum_{l=1}^{N}(\partial_{a}T_{rl}(\bm{k}))T^{\dagger}_{ls}(\bm{k})$ and $T^{\dagger}_{ls}(\bm{k})\equiv\big{(}T^{\dagger}(\bm{k})\big{)}_{ls}$. Note that since $T(\bm{k})$ is block-diagonal, so too is $\mathcal{T}^{a}(\bm{k})$. Thus, $(\xi^{\mathfrak{v}})^{a}_{vc}=\sum_{v^{\prime}c^{\prime}}T^{\dagger}_{v,v^{\prime}}(\xi^{\mathfrak{u}})^{a}_{v^{\prime}c^{\prime}}T_{c^{\prime}c}$. Plugging this into Eq. (S50) we find
$\bar{\alpha}_{\mathfrak{v}}^{li}$ $=\frac{e^{2}}{2\hbar c}\epsilon^{iab}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\text{Re}\Bigg{[}-i\sum_{vv^{\prime}c}(\xi^{\mathfrak{u}})^{a}_{cv^{\prime}}\big{(}(\xi^{\mathfrak{u}})^{b}_{v^{\prime}v}+\mathcal{T}^{b}_{v^{\prime}v}\big{)}(\xi^{\mathfrak{u}})^{l}_{vc}+i\sum_{vcc^{\prime}}\big{(}(\xi^{\mathfrak{u}})^{a}_{cc^{\prime}}+\mathcal{T}^{a}_{cc^{\prime}}\big{)}(\xi^{\mathfrak{u}})^{b}_{c^{\prime}v}(\xi^{\mathfrak{u}})^{l}_{vc}\Bigg{]}$
$=\bar{\alpha}_{\mathfrak{u}}^{li}+\frac{e^{2}}{2\hbar c}\epsilon^{iab}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\text{Re}\Bigg{[}-i\sum_{vv^{\prime}c}(\xi^{\mathfrak{u}})^{a}_{cv^{\prime}}\mathcal{T}^{b}_{v^{\prime}v}(\xi^{\mathfrak{u}})^{l}_{vc}+i\sum_{vcc^{\prime}}\mathcal{T}^{a}_{cc^{\prime}}(\xi^{\mathfrak{u}})^{b}_{c^{\prime}v}(\xi^{\mathfrak{u}})^{l}_{vc}\Bigg{]}.$ (S53)
That is, calculations of $\bar{\alpha}^{li}_{\mathfrak{v}}$ can always be made using the components of the Berry connection $\xi^{\mathfrak{u}}$ induced by some reference gauge $\mathfrak{u}$ and the components of the $\mathcal{T}^{a}$ obtained from the unitary transformation $T$ that relates $\mathfrak{v}$ and $\mathfrak{u}$.
S2.3.2 Specific considerations for our model
In past work [15] we studied the non-magnetic analog of the model employed here and considered whether there exists a gauge $\mathfrak{v}$ in which $\bar{\alpha}_{\mathfrak{v}}^{li}$ vanishes (independent of the choice of model parameters). That model has $N=4$ and we considered the initial electronic state of the material to be such that $M=2$ (i.e., $v\in\{1,2\}$ and $c\in\{3,4\}$). Consider general smooth global (and periodic) gauges $\mathfrak{u}$ and $\mathfrak{v}$ in the total Bloch bundle, the components of which are related pointwise as in Eq. (S51). We only consider gauges in the total Bloch bundle that decompose as a product of smooth gauges on the occupied and unoccupied subbundles thereof. For gauges of this type, in the non-magnetic model gauge transformations are represented by a $4\times 4$ block-diagonal unitary matrix consisting of two $2\times 2$ unitary blocks. The question is whether or not one can choose the two $2\times 2$ blocks of $T$ in such a way that $\mathcal{T}_{vv^{\prime}}^{a}$ and $\mathcal{T}_{cc^{\prime}}^{a}$ are such that the integrand of Eq. (S53) vanishes. (The integral as a whole may accidentally vanish in some gauge for a given set of material parameters due to fine tuning, but if the Hamiltonian parameters are slightly altered this fine tuning cancellation would be destroyed – it would be more general if the integrand itself always vanished.) Indeed in past work [15] we explicitly found a gauge in which this occurred.
Let’s now focus on the anti-ferromagnetic model. We first employ our particular gauge choice $\mathfrak{u}$ defined by the Bloch energy eigenvectors in Eq. (S40,S41). Considering the $i=l$ component of Eq. (S50) and using the relations (S44,S45), we find
$\bar{\alpha}^{ll}_{\mathfrak{u}}$ $=\frac{e^{2}}{2\hbar c}\epsilon^{lab}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\text{Re}\Bigg{[}\sum_{v=1}^{2}\sum_{c=7}^{8}\Bigg{(}-i\sum_{v^{\prime}=1}^{2}(\xi^{\mathfrak{u}})^{a}_{cv^{\prime}}(\xi^{\mathfrak{u}})^{b}_{v^{\prime}v}+i\sum_{c^{\prime}=7}^{8}(\xi^{\mathfrak{u}})^{a}_{cc^{\prime}}(\xi^{\mathfrak{u}})^{b}_{c^{\prime}v}\Bigg{)}(\xi^{\mathfrak{u}})^{l}_{vc}$
$\qquad\qquad\qquad\qquad\qquad\quad+\sum_{v=3}^{4}\sum_{c=5}^{6}\Bigg{(}-i\sum_{v^{\prime}=3}^{4}(\xi^{\mathfrak{u}})^{a}_{cv^{\prime}}(\xi^{\mathfrak{u}})^{b}_{v^{\prime}v}+i\sum_{c^{\prime}=5}^{6}(\xi^{\mathfrak{u}})^{a}_{cc^{\prime}}(\xi^{\mathfrak{u}})^{b}_{c^{\prime}v}\Bigg{)}(\xi^{\mathfrak{u}})^{l}_{vc}\Bigg{]}.$ (S54)
In order for Eq. (S54) to generally vanish over the parameter space of the model, either the terms in each of the ()($\ldots$)( … ) must cancel one another or the two ()($\ldots$)( … ) terms must cancel each other. Since there is generally no relation between, for example, $(\xi^{\mathfrak{u}})^{l}_{17}$ and $(\xi^{\mathfrak{u}})^{l}_{35}$, the ()($\ldots$)( … ) terms cannot generally cancel each other, thus the second possibility is eliminated. The first possibility does not occur either, since there is generally no relationship between $(\xi^{\mathfrak{u}})^{a}_{12}$ and $(\xi^{\mathfrak{u}})^{a}_{78}$ in the first ()($\ldots$)( … ) of Eq. (S54) nor between $(\xi^{\mathfrak{u}})^{a}_{34}$ and $(\xi^{\mathfrak{u}})^{a}_{56}$ in the second ()($\ldots$)( … ). (In the original gauge $\mathfrak{u}$ employed in the non-magnetic case there was such a relationship between $(\xi^{\mathfrak{u}})^{a}_{v^{\prime}v}$ and $(\xi^{\mathfrak{u}})^{a}_{cc^{\prime}}$ for $v\neq v^{\prime}$ and $c\neq c^{\prime}$.)
What about in some other smooth global Hamiltonian gauge $\mathfrak{v}$? Due to the degeneracy in the model we employ, at each $\bm{k}\in\text{BZ}$ the transformation $T(\bm{k})$ is represented by a block-diagonal matrix consisting of four $2\times 2$ unitary blocks. Then
$\bar{\alpha}^{ll}_{\mathfrak{v}}$ $=\frac{e^{2}}{2\hbar c}\epsilon^{lab}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\text{Re}\Bigg{[}\sum_{v=1}^{2}\sum_{c=7}^{8}\Bigg{(}-i\sum_{v^{\prime}=1}^{2}(\xi^{\mathfrak{u}})^{a}_{cv^{\prime}}\big{(}(\xi^{\mathfrak{u}})^{b}_{v^{\prime}v}+\mathcal{T}^{b}_{v^{\prime}v}\big{)}+i\sum_{c^{\prime}=7}^{8}\big{(}(\xi^{\mathfrak{u}})^{a}_{cc^{\prime}}+\mathcal{T}^{a}_{cc^{\prime}}\big{)}(\xi^{\mathfrak{u}})^{b}_{c^{\prime}v}\Bigg{)}(\xi^{\mathfrak{u}})^{l}_{vc}$
$\qquad\qquad\qquad\qquad\qquad\quad+\sum_{v=3}^{4}\sum_{c=5}^{6}\Bigg{(}-i\sum_{v^{\prime}=3}^{4}(\xi^{\mathfrak{u}})^{a}_{cv^{\prime}}\big{(}(\xi^{\mathfrak{u}})^{b}_{v^{\prime}v}+\mathcal{T}^{b}_{v^{\prime}v}\big{)}+i\sum_{c^{\prime}=5}^{6}\big{(}(\xi^{\mathfrak{u}})^{a}_{cc^{\prime}}+\mathcal{T}^{a}_{cc^{\prime}}\big{)}(\xi^{\mathfrak{u}})^{b}_{c^{\prime}v}\Bigg{)}(\xi^{\mathfrak{u}})^{l}_{vc}\Bigg{]}.$ (S55)
Again, the lack of a relation between, for example, $(\xi^{\mathfrak{u}})^{l}_{17}$ and $(\xi^{\mathfrak{u}})^{l}_{35}$, means that the two ()($\ldots$)( … ) terms cannot generally cancel one another. So for the integrand to vanish there must be cancellation within each ()($\ldots$)( … ). Using the relations $(\xi^{\mathfrak{u}})^{a}_{17}=-(\xi^{\mathfrak{u}})^{a}_{82}$ and $(\xi^{\mathfrak{u}})^{a}_{18}=(\xi^{\mathfrak{u}})^{a}_{72}$ given in (S44), the first ()($\ldots$)( … ) term in Eq. (S55) can be rewritten as
$\epsilon^{lab}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\text{Re}\Bigg{[}\sum_{v=1}^{2}\sum_{c=7}^{8}\Bigg{(}-i\sum_{v^{\prime}=1}^{2}(\xi^{\mathfrak{u}})^{a}_{cv^{\prime}}\big{(}(\xi^{\mathfrak{u}})^{b}_{v^{\prime}v}+\mathcal{T}^{b}_{v^{\prime}v}\big{)}+i\sum_{c^{\prime}=7}^{8}\big{(}(\xi^{\mathfrak{u}})^{a}_{cc^{\prime}}+\mathcal{T}^{a}_{cc^{\prime}}\big{)}(\xi^{\mathfrak{u}})^{b}_{c^{\prime}v}\Bigg{)}(\xi^{\mathfrak{u}})^{l}_{vc}\Bigg{]}$
$=\epsilon^{lab}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\text{Re}\Bigg{[}i(\xi^{\mathfrak{u}})^{l}_{17}\bigg{(}\Big{\{}(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{11}+(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{77}-(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{22}-(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{88}\Big{\}}(\xi^{\mathfrak{u}})^{b}_{71}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+2(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{78}(\xi^{\mathfrak{u}})^{b}_{81}+2(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{12}(\xi^{\mathfrak{u}})^{b}_{18}\bigg{)}$
$\quad\qquad\qquad\qquad\qquad+i(\xi^{\mathfrak{u}})^{l}_{18}\bigg{(}\Big{\{}(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{11}+(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{88}-(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{22}-(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{77}\Big{\}}(\xi^{\mathfrak{u}})^{b}_{81}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+2(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{87}(\xi^{\mathfrak{u}})^{b}_{71}-2(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{21}(\xi^{\mathfrak{u}})^{b}_{17}\bigg{)}\Bigg{]},$ (S56)
[…]
Using the relations $(\xi^{\mathfrak{u}})^{a}_{35}=-(\xi^{\mathfrak{u}})^{a}_{64}$ and $(\xi^{\mathfrak{u}})^{a}_{36}=(\xi^{\mathfrak{u}})^{a}_{54}$, and $(\xi^{\mathfrak{u}})^{a}_{33}=-(\xi^{\mathfrak{u}})^{a}_{44}$ and $(\xi^{\mathfrak{u}})^{a}_{55}=-(\xi^{\mathfrak{u}})^{a}_{66}$ given in (S44), analogous arguments apply to the the second bracketed term in Eq. (S55). The sum total of these arguments is that there exists a smooth global Hamiltonian gauge $\mathfrak{v}$ in which $\bar{\alpha}^{ll}_{\mathfrak{v}}$ vanishes independent of material parameters only if there exists a $T$ relating smooth Hamiltonian gauges $\mathfrak{u}$ and $\mathfrak{v}$ that satisfies $(\xi^{\mathfrak{u}}+\mathcal{T})^{a}_{nm}=\delta_{nm}\partial_{a}\phi(\bm{k})$ (thus from Eq. (S52) we have $(\xi^{\mathfrak{v}})^{a}_{nm}=\delta_{nm}\partial_{a}\phi(\bm{k})$) for $n$ and $m$ in the same degenerate $2\times 2$ block of bands. And since $\mathfrak{v}$ is a Hamiltonian gauge (i.e., $T(\bm{k})$ and $\mathcal{T}^{a}(\bm{k})$ are block diagonal consisting of four $2\times 2$ blocks), Eq. (S52) implies that when $n$ and $m$ are not in the same degenerate $2\times 2$ block of bands, $(\xi^{\mathfrak{v}})^{a}_{nm}$ satisfies an analogous version of Eq. (S45); that is, $(\xi^{\mathfrak{v}})^{a}_{13},\ldots,(\xi^{\mathfrak{v}})^{a}_{86}$ vanishes unless $a=z$. If we assume that such a $T$ (and therefore $\mathfrak{v}$ related to $\mathfrak{u}$ by that $T$) exists, then
$\alpha^{\mathfrak{v}}_{\text{CS}}$ $=-\frac{e^{2}}{2\hbar c}\epsilon^{abd}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\left(\sum_{vv^{\prime}}(\xi^{\mathfrak{v}})^{a}_{vv^{\prime}}\partial_{b}(\xi^{\mathfrak{v}})^{d}_{v^{\prime}v}-\frac{2i}{3}\sum_{vv^{\prime}v_{1}}(\xi^{\mathfrak{v}})^{a}_{vv^{\prime}}(\xi^{\mathfrak{v}})^{b}_{v^{\prime}v_{1}}(\xi^{\mathfrak{v}})^{d}_{v_{1}v}\right)=0,$ (S57)
where we have used $\epsilon^{abd}\partial_{b}\partial_{d}\phi=0$ and $\epsilon^{abd}(\partial_{b}\phi)(\partial_{d}\phi)=0$. This result is independent of material parameters. Indeed this is a contradiction since we have explicitly shown above that this model supports regions of parameter space in which the electronic ground state (at half filling) is $\mathbb{Z}_{2}$-odd and $\alpha^{\mathfrak{v}}_{\text{CS}}\text{ mod }\frac{e^{2}}{hc}=\frac{e^{2}}{2hc}$. Thus, there does not exist a smooth global Hamiltonian gauge $\mathfrak{v}$ in which $\bar{\alpha}^{ll}_{\mathfrak{v}}$ vanishes independent of material parameters. That is, in the bulk 3D AFM tight-binding model considered here, the magnetoelectric response cannot always be made entirely itinerant (with respect to a smooth global and periodic Hamiltonian gauge), in contrast to the non-magnetic case [15].